### v1

# Optimal Piecewise-based Mechanism for Collecting Bounded Numerical Data under Local Differential Privacy

Abstract.
Numerical data with bounded domains is a common data type in personal devices, such as wearable sensors. While the collection of such data is essential for third-party platforms, it raises significant privacy concerns. Local differential privacy (LDP) has been shown as a framework providing provable individual privacy, even when the third-party platform is untrusted. For numerical data with bounded domains, existing state-of-the-art LDP mechanisms are piecewise-based mechanisms, which are not optimal, leading to reduced data utility.
This paper investigates the optimal design of piecewise-based mechanisms to maximize data utility under LDP. We demonstrate that existing piecewise-based mechanisms are heuristic forms of the $3$-piecewise mechanism, which is far from enough to study optimality. We generalize the $3$-piecewise mechanism to its most general form, i.e. $m$-piecewise mechanism with no pre-defined form of each piece. Under this form, we derive the closed-form optimal mechanism by combining analytical proofs and off-the-shelf optimization solvers. Next, we extend the generalized piecewise-based mechanism to the circular domain (along with the classical domain), defined on a cyclic range where the distance between the two endpoints is zero. By incorporating this property, we design the optimal mechanism for the circular domain, achieving significantly improved data utility compared with existing mechanisms.
Our proposed mechanisms guarantee optimal data utility under LDP among all generalized piecewise-based mechanisms. We show that they also achieve optimal data utility in two common applications of LDP: distribution estimation and mean estimation. Theoretical analyses and experimental evaluations prove and validate the data utility advantages of our proposed mechanisms.
local differential privacy, numerical data privacy, bounded domain, circular data
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
Numerical data with bounded domains is a fundamental data type in personal devices. These bounded domains can be categorized into two types: linear ranges, such as sensor readings in $[0,1)$, referred to as classical domain; and cyclic ranges, such as angular measurements in $[0,2\pi)$, referred to as circular domain. These types of data are crucial for third-party platforms to provide personalized services. However, collecting such data involves particular privacy concerns, as the third-party collectors are potentially untrusted and the data often contain sensitive information. Simple anonymization techniques (Yang et al., 2013; Niu et al., 2014) have been proven insufficient to prevent privacy leakage (Wikipedia contributors, 2023; Narayanan and Shmatikov, 2008; Yu et al., 2007). Therefore, a provable privacy guarantee is necessary when collecting these sensitive data.
Local differential privacy (LDP) serves as a de facto standard, providing an input-independent formal guarantee regarding the difficulty of inferring sensitive data. Through the LDP mechanism, sensitive data are randomly perturbed before being sent to an untrusted collector. The randomization ensures a sufficient level of indistinguishability (indicated by the privacy parameter $\varepsilon$). Consequently, any observation over the randomized data is essentially powerless to infer the sensitive data. Laplace Mechanism (Dwork et al., 2006) is a classical LDP mechanism for numerical data privacy. It adds random noise, drawn from a Laplace distribution determined by $\varepsilon$, to the sensitive data. However, the unbounded noise of the Laplace mechanism makes it unsuitable for bounded domains.
State-of-the-art LDP mechanisms for numerical data with bounded domains are piecewise-based mechanisms (Wang et al., 2019; Li et al., 2020; Ma et al., 2024). They are widely used as building blocks to provide provable privacy guarantees in various scenarios, such as in sensors networks and federated learning. Piecewise-based mechanisms randomize the sensitive data to a value sampled from a carefully designed piecewise probability distribution. Existing instantiations use different pieces and probabilities, but all are designed for classical domains. Their applicability to other types of bounded domains, e.g. circular domain of angular sensors that commonly appear in personal devices, is unexplored.
Data utility is the most crucial metric for LDP mechanisms, typically measured by the distance between the randomized data and the sensitive data. While the privacy level is theoretically guaranteed by the privacy parameter $\varepsilon$, a mechanism with better data utility allows for more accurate analysis. The data utility of a piecewise-based mechanism is determined by the distance metric, pieces, and their respective probabilities. Unfortunately, none of the existing instantiations of piecewise-based mechanisms are shown to be optimal in terms of data utility, indicating potential for improving analysis accuracy without compromising privacy. These situations highlight the need for optimal piecewise-based mechanisms.
The optimality of piecewise-based mechanisms remains a challenging problem. We will see that existing instantiations are heuristic forms of the $3$-piecewise mechanism (TPM). As a special case with $3$ pieces and pre-defined forms of pieces, TPM is far from enough to study the optimality of piecewise-based mechanisms.  From the evidence of the staircase Laplace mechanism (Geng and Viswanath, 2014) for unbounded numerical data, the asymptotically optimal mechanism has a staircase (multiple pieces) form. For categorical data, a staircase Randomized Response mechanism (SRR) (Wang et al., 2022) improves data utility in location collection compared to the general RR mechanism. Numerous indicators suggest that increasing the variety of probabilities in the data domain, i.e. using more pieces, can improve data utility.  In light of these examples, a fundamental question for piecewise-based mechanisms is: what is the optimal instantiation of piecewise-based mechanism? In the design of piecewise-based mechanisms, the number of pieces, their probabilities and sizes can be arbitrary. Finding the optimal instantiation within such a large design space is challenging, as it requires optimizing the number of pieces, their probabilities and sizes simultaneously.
This paper studies the optimality of piecewise-based mechanism in its most general form.111 This means that we consider all possible forms of piecewise distributions on a bounded domain, ensuring the most comprehensive generalization. We extend TPM into a generalized piecewise-based mechanism (GPM) that is $m$-piece, with each piece having no pre-defined form. Under GPM, we formulate an optimization problem to minimize the distance between the sensitive and randomized data. By combining the solutions of the optimization problem with analytical proofs, we derive the closed-form optimal GPM for classical domains.  For circular domains, where distance metrics have a unique property (e.g. the distance between 00 and $2\pi$ is zero), we incorporate this property into mechanism design and link the solving of the optimal mechanism to problems in the classical domain. Table 1 summarizes the key features of this paper in comparison with existing works. Particularly, our contributions are as follows:
  • (Solving framework) To the best of our knowledge, this is the first work to study the closed-form optimal piecewise-based mechanism under its most general form. We propose a framework that integrates analytical proofs with off-the-shelf optimization solvers to derive the closed-form optimal mechanism. This approach establishes a feasible foundation for achieving optimal data utility under LDP for numerical data with bounded domains.
  • (Closed-form instantiations) We provide closed-form optimal mechanisms for the classical domain and the circular domain. As alternatives to existing mechanisms, they can be directly used as building blocks in sensor networks and federated learning, etc, while guaranteeing optimal data utility among all piecewise-based mechanisms under LDP.
  • (Theoretical and experimental evaluations) We provide theoretical analyses of data utility and experimental evaluations on two common applications of LDP: distribution and mean estimation. The results prove and validate our mechanisms’ advantages over existing mechanisms. Codes are available at https://github.com/ZhengYeah/Optimal-GPM.
Structure. The main part of this paper is organized as follows: After the preliminaries, we present the optimal piecewise-based mechanism for the classical domain in Section 3. Section 4 focuses on the circular domain and derives its optimal mechanism. Following this, Section 5 discusses the optimality when applying to two common tasks: distribution and mean estimation.
2. Preliminaries
This section formulates the problem and the concept of local differential privacy (LDP). We present existing instantiations of TPM and their limitations, which motivate our proposed optimal generalized piecewise-based mechanism (OGPM).
2.1. Problem Formulation
We consider a typical data collection schema that consists of a set of users and one collector. Each user has a numerical sensitive data $x_{i}\in\mathcal{D}$, where $\mathcal{D}$ is a continuous and bounded domain. The collector needs to collect data from users for statistical estimations, such as the mean value and distribution of the data.
However, the collector is untrusted and may attempt to infer users’ sensitive data. To protect privacy, each user locally randomizes their sensitive data using a privacy mechanism $\mathcal{M}:\mathcal{D}\to\tilde{\mathcal{D}}$, then sends $y_{i}=\mathcal{M}(x_{i})$ to the collector.
We seek to design an optimal $\mathcal{M}$ that maximizes the data utility by minimizing the distance between the sensitive data $x_{i}$ and the reported data $y_{i}$, while ensuring $\varepsilon$-LDP (Definition 2.1).
2.2. Local Differential Privacy
Definition 2.1 ($\varepsilon$-LDP (Duchi et al., 2016)).
A randomization mechanism $\mathcal{M}:\mathcal{D}\to\tilde{\mathcal{D}}$ satisfies $\varepsilon$-LDP, if for two arbitrary inputs $x_{1}$ and $x_{2}$, the probability ratio of outputting the same $y$ is bounded:
$\forall x_{1},x_{2}\in\mathcal{D},\forall y\in\tilde{\mathcal{D}}:\frac{\Pr[\mathcal{M}(x_{1})=y]}{\Pr[\mathcal{M}(x_{2})=y]}\leq\exp(\varepsilon).$
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by probability density function ($pdf$). Intuitively, $\varepsilon$-LDP represents the difficulty of distinguishing between $x_{1}$ and $x_{2}$ given $y$. A lower privacy parameter $\varepsilon\in[0,+\infty)$ means higher privacy.  For example, $\varepsilon=0$ requires $\mathcal{M}$ maps two arbitrary inputs to any output $y$ with the same probability, thus $y$ contains no distribution information about $x$, making any hypothesis-testing method to infer the sensitive $x$ powerless even with known $\mathcal{M}$.
2.3. Piecewise-based Mechanisms
Table 1. OGPM vs existing instantiations of TPM.
Domain Optimality Closed form Estimation
PM (Wang et al., 2019) Classical No Yes Mean
SW (Li et al., 2020) No Yes Distribution
PTT (Ma et al., 2024) Partly* No Mean
This paper (OGPM) Classical & circular Yes Yes Mean & distribution
  • * Proved the existence of the optimal under TPM, but did not give closed-form instantiations. Appendix B.1 provides detailed discussion.
When the input domain $\mathcal{D}$ is both continuous and bounded, the state-of-the-art mechanisms to achieve $\varepsilon$-LDP are piecewise-based mechanisms. We summarize these mechanisms as heuristic forms of the following definition.
Definition 2.2.
$3$-piecewise mechanism (TPM) $\mathcal{M}:\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$ is a family of probability density functions that, given input $x\in\mathcal{D}$, outputs $y\in\tilde{\mathcal{D}}$ according to
$pdf[\mathcal{M}(x)=y]=\begin{cases}p_{\varepsilon}&\text{if}\ y\in[l_{x,\varepsilon},r_{x,\varepsilon}],\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{if }y\in\tilde{\mathcal{D}}_{\varepsilon}\,\backslash\,[l_{x,\varepsilon},r_{x,\varepsilon}],\end{cases}$
where $p_{\varepsilon}$ is a variable determined solely by $\varepsilon$,222 Otherwise, if $p_{\varepsilon}$ varies with $x$, it violates the $\varepsilon$-LDP constraint because the probability ratio outputting the same $y$ from $x_{1}$ and $x_{2}$ is not bounded by $\exp(\varepsilon)$. while $l_{x,\varepsilon}$ and $r_{x,\varepsilon}$ depend on $x$ and $\varepsilon$. The output domain $\tilde{\mathcal{D}}_{\varepsilon}\supset\mathcal{D}$ is an enlarged domain depending on $\varepsilon$.
TPM samples the output $y$ for each $x$ from a piecewise distribution. This sampling is with a higher probability $p_{\varepsilon}$ within $[l_{x,\varepsilon},r_{x,\varepsilon}]$ and a lower probability $p_{\varepsilon}/\exp(\varepsilon)$ within the remaining two pieces $\tilde{\mathcal{D}}\,\backslash\,[l_{x,\varepsilon},r_{x,\varepsilon}]$, satisfying the $\varepsilon$-LDP constraint.
Instantiations. In TPM, the parameters are the central interval $[l_{x,\varepsilon},r_{x,\varepsilon}]$, its probability $p_{\varepsilon}$, and the output domain $\tilde{\mathcal{D}}_{\varepsilon}$. Different instantiations of those parameters yield different existing mechanisms (Wang et al., 2019; Li et al., 2020; Ma et al., 2024). For example, PM (Wang et al., 2019) is the first instantiation of TPM, defined on $[-1,1]\to[-C_{\varepsilon},C_{\varepsilon}]$, where $C_{\varepsilon}$ is a variable determined solely by $\varepsilon$ and the central interval has a fixed length $r_{x,\varepsilon}-l_{x,\varepsilon}=C_{\varepsilon}-1$.
Data utility metric. To quantify the data utility of different instantiations, we consider the general $L_{p}$-similar error metric (i.e. $|y-x|^{p}$) as a loss function $\mathcal{L}:\mathbb{R}\to\mathbb{R}$. Thus, the error is:
(1) $Err(x)=\int_{\tilde{\mathcal{D}}}\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}\mathrm{d}y,$
where $\mathcal{P}_{\mathcal{M}(x)}$ is the pdf defined by $\mathcal{M}(x)$. $Err(x)$ illustrates the expected error when applying $\mathcal{M}$ on $x$ under the loss function $\mathcal{L}$. For example, $\mathcal{L}(y,x)\coloneq|y-x|$ is the absolute error, and $\mathcal{L}(y,x)\coloneq(y-x)^{2}$ is the square error. Then $Err(x)$ corresponds to the mean absolute error (MAE) and mean square error (MSE) (Wang et al., 2019; Ma et al., 2024), respectively. Lower $Err(x)$ indicates better data utility.
Limitations of TPM. Existing instantiations of TPM have the following limitations.
  • Not optimal in data utility. None of the existing instantiations provided closed forms for the optimal data utility. Meanwhile, they also assume invariable length $r_{x,\varepsilon}-l_{x,\varepsilon}$ of the central piece for all $x$, and symmetric probability $p_{\varepsilon}/\exp(\varepsilon)$ for the remaining two pieces. However, a general-form piecewise-based mechanism can have more pieces, unfixed piece lengths, and asymmetric probabilities, potentially improving data utility.
  • Limited applicability. Existing instantiations of TPM have enlarged and unfixed output domains $\tilde{\mathcal{D}}_{\varepsilon}\supset\mathcal{D}$. Enlarged output domain incurs applicability issues in scenarios where the collector requires the output domain to align with the input domain (i.e. $\tilde{\mathcal{D}}=\mathcal{D}$),333 While post-processing the output by truncating it to $\mathcal{D}$ is possible, this approach may still result in low data utility. Sections 7.1.2 and 7.1.3 provide comparisons with mechanisms that include truncation. such as in common sensor-based services.
3. Generalized Piecewise-based Mechanism
This section generalizes TPM to its most general form (GPM). We introduce a framework for deriving the closed-form optimal GPM for the classical domain.
Definition 3.1.
Generalized $m$-piecewise mechanism ($m$-GPM) $\mathcal{M}:\mathcal{D}\to\tilde{\mathcal{D}}$ is a family of probability density functions that, given input $x\in\mathcal{D}$, outputs $y\in\tilde{\mathcal{D}}$ according to
$pdf[\mathcal{M}(x)=y]$ =$=\begin{dcases}p_{1,\varepsilon}&\text{if}\ y\in[l_{1,x,\varepsilon},r_{1,x,\varepsilon}),\\ \ \vdots&\vdots\quad\quad\ \ \vdots\\ p_{m,\varepsilon}&\text{if}\ y\in[l_{m,x,\varepsilon},r_{m,x,\varepsilon}),\\ \end{dcases}$
$\forall i,j\in[m],$ $\max\frac{p_{i,\varepsilon}}{p_{j,\varepsilon}}\leq\exp(\varepsilon),$
where $[m]\coloneq\{1,...,m\}$. Each probability $p_{i,\varepsilon}$ depends solely on privacy parameter $\varepsilon$, while interval boundaries $l_{i,x,\varepsilon}$ and $r_{i,x,\varepsilon}$ depend on both $x$ and $\varepsilon$.
An $m$-GPM partitions its output domain into $m$ pieces, assigning probability $p_{i,\varepsilon}$ to each piece $[l_{i,x,\varepsilon},r_{i,x,\varepsilon})$. The probabilities $p_{i,\varepsilon}$ are independent of the input $x$, and their ratios must be bounded by $\exp(\varepsilon)$ to satisfy $\varepsilon$-LDP. For notational clarity, we omit subscripts $x$ and $\varepsilon$ when their context is clear. Additionally, $\mathcal{M}$ must satisfy standard probability requirements: non-negativity ($p_{i}\geq 0$), continuity ($r_{i}=l_{i+1}$), and normalization. TPM is a special case of GPM with $m=3$.
Finding the optimal GPM requires determining both the optimal number of pieces $m$ and the corresponding $p_{i,\varepsilon},l_{i,x,\varepsilon},r_{i,x,\varepsilon}$. Due to the infinite possibilities for $m\in\mathbb{N}^{+}$ and the resulting $3m$ variables, analytical solutions are computationally intractable. We therefore propose a framework that combines analytical proofs with off-the-shelf optimization solvers.
3.1. Framework for Deriving the Optimal GPM
To derive the closed-form optimal GPM, we (i) formulate finding the optimal $m$-GPM as an optimization problem; (ii) determine the optimal $m$ based on the solutions of the optimization problem; (iii) derive the optimal closed-form expression (among all $m$-GPM).
Optimal $m$-GPM. To find the optimal GPM instantiation with $m$ pieces, we need to determine the variables $p_{i}$, $l_{i}$, and $r_{i}$. Any feasible assignment of these variables yields a mechanism $\mathcal{M}$ whose utility can be measured by $Err(x)$ from Formula (1). Finding the optimal $m$-GPM requires solving a min-max optimization problem that minimizes the worst-case error over all possible inputs $x$:444 Worst-case error is the most common utility metric in mechanism design (Wang et al., 2019; Ma et al., 2024). We can also optimize the error at other specific points, see Section 6.
[…]
This formulation yields the optimal $x$-independent $p_{i}$ values and the corresponding $l_{i,x},r_{i,x}$ for the worst-case input $x$. However, since these $l_{i,x},r_{i,x}$ may not be optimal for other inputs, we need a second optimization step using the obtained optimal $p_{i}$:
[…]
Together, these two steps determine the optimal instantiation of $m$-GPM for any given domain mapping $\mathcal{D}\to\tilde{\mathcal{D}}$, distance metric $\mathcal{L}$, piece number $m$, privacy parameter $\varepsilon$, and input $x$. Figure 1 illustrates this solving process.
Figure 1. Solving flow for the optimal $m$-GPM. Two arrows indicate problems in (2) and (3).
Challenges. Nonetheless, solving Formulation (2) has practical difficulties. Even if it can be solved, there is still a gap between the solved optimal $m$-GPM and the closed-form optimal GPM (among all $m$-GPM). We detail them as follows.
  • (Solving difficulty) Formulation (2) is a min-max problem, and the integrand $\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}$ is non-linear; it is a non-convex problem whose global optimal hard to solve.
  • (Optimal $m$) The solved optimal is only for $m$-GPM given $m$. It is necessary to find the optimal piece number $m$.
  • (Closed form) For practical usage, we need closed-form $p_{i}$, $l_{i,x}$ and $r_{i,x}$ (i.e. their relationships with $\varepsilon$ and $x$), rather than their specific values for every $\varepsilon$ and $x$.
Solutions. We address the challenges and gaps with the following solutions.
  • Formulation (2) can be simplified to two bilinear optimization problems that can be solved by off-the-shelf solvers. (Section 3.2.1)
  • If the optimal ($m+1$)-GPM is identical to the $m$-GPM, then $m$ is the optimal piece number. (Section 3.2.2)
  • Leveraging the above results, the optimal closed-form $p_{i}$, $l_{i,x}$ and $r_{i,x}$ can be obtained by analytical deduction (for TPM) or numerical regression (for any $m$-GPM). (Section 3.2.3)
Following these solutions, we can obtain the closed-form optimal GPM (among all $m$) for any $\varepsilon$ given $\mathcal{D}\to\tilde{\mathcal{D}}$ and $\mathcal{L}$. Before presenting the detailed solutions, we first discuss the conditions under which the obtained GPM is optimal.
Conditions for optimality. When discussing optimality, the following aspects should be specified: (i) the error metric, (ii) the data domain and family of mechanisms, (iii) the strength of the optimality, and (iv) whether post-processing is allowed. In this paper, the optimality of GPM is defined with respect to: (i) the worst-case $L_{p}$-similar error metric, (ii) bounded numerical domains $\mathcal{D}\to\tilde{\mathcal{D}}$ and mechanisms based on piecewise distributions, (iii) minimization of error value (not asymptotic or order-of-magnitude optimality), and (iv) without post-processing. These conditions are widely applicable in practice and literature. However, varying any of them may lead to different optimality results. Appendix B.2 provides a detailed discussion of these conditions and related optimalities.
3.2. Detailed Solutions
3.2.1. Solution 1: Simplified Form
The min-max problem in Formulation (2) can be simplified. The key observation is that its inner maximization term, $\max_{x}$, has a closed form, i.e. the worst-case error is from the endpoints of $\mathcal{D}$. Lemma 3.2 states this observation.
Lemma 3.2.
Assume $\mathcal{D}=[a,b)$, the objective of Formulation (2) can be simplified to
$\min_{p_{i},l_{i,x},r_{i,x}}\max_{x\in\{a,b\}}\int_{\tilde{\mathcal{D}}}\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}\mathrm{d}y.$
Proof.
(Sketch) The key of the proof is that each integral on $\tilde{\mathcal{D}}$ is convex function w.r.t $x$. Thus, their non-negative weighted sum is also convex. According to the Bauer maximum principle (Wikipedia, 2024), the maximum is achieved at the endpoints of $\mathcal{D}$, i.e. $x=a$ or $b$555 We use $[a,b)$ to denote the domain for consistency with the form of pieces in GPM. This is the same as $[a,b]$ in implementation. Appendix A.1 provides the full proof. ∎
Complexity. (i) Lemma 3.2 simplifies Formulation (2) to two bilinear optimization problems, i.e. when $x=a$ and $x=b$ respectively. The integrand $\mathcal{L}(y,a)\mathcal{P}_{\mathcal{M}(a)}$ includes terms such as $p_{i}l_{i}$ and $p_{i}r_{i}$, which involve multiplications of two variables. This problem can be solved by off-the-shelf solvers such as Gurobi (noa, 2024), which employs stochastic Branch and Bound (Gupte et al., 2017) method to handle the bilinear terms.  (ii) The number of variables is at most $3m$, which can be efficiently solved for small $m$. For example, we can obtain the exact optimal for $m\leq 7$ and $\mathcal{L}=|y-x|$ within $2$ seconds. Furthermore, solvers generally provide over- and under-approximation for bilinear problems (Nekst-Online, 2016). We can obtain solutions with $\leq 1\%$ gap from the optimal for $m\leq 19$ within $1$ minute.  (iii) Formulation (3) for solving $l_{i}$ and $r_{i}$ has at most $2m$ variables and without $p_{i}$ terms, it is a toy-size problem when $m$ is small.
Encoding details. We need to encode the problem in Lemma 3.2 to simple mathematical expressions that can be handled by the solver. If we instantiate $\mathcal{L}=|y-x|$ and focus on the left endpoint $x=a$, this problem becomes:
$\begin{split}&\min_{p_{i},l_{i},r_{i}}\sum_{i=1}^{m}p_{i}\int_{l_{i}}^{r_{i}}(y-a)\mathrm{d}y\\ =&\min_{p_{i},l_{i},r_{i}}\sum_{i=1}^{m}\frac{p_{i}}{2}\left((r_{i}-a)^{2}-(l_{i}-a)^{2}\right).\end{split}$
This problem is a bilinear optimization problem. The highest-degree term is $p_{i}\cdot r_{i}\cdot r_{i}$, which can be reformulated as $p_{i}\cdot t$ with $t=r_{i}\cdot r_{i}$, i.e. multiplication of bilinear terms. Such problems can be solved by off-the-shelf bilinear solvers (noa, 2024). The other problem in Formulation (3) can be encoded similarly but is much easier due to constant $p_{i}$.
3.2.2. Solution 2: Optimal Piece Number
Although we can obtain the optimal $m$-GPM for any $m$ given sufficient time, solving for each $m$ is unnecessary. The following lemma provides a theoretical basis for capping the optimal number of pieces.
Lemma 3.3.
For all possible $\varepsilon$ and $x$, if the optimal ($m+1$)-GPM is the same as the $m$-GPM, then the optimal piece number is $m$.
Proof.
(Sketch) The key insight is that, if the optimal piece number is not $m$, i.e. an additional piece can lower the error, then this additional piece will be captured by the optimal ($m+1$)-GPM.  Therefore, (i) if $m$ is not the optimal piece number, then the optimal ($m+1$)-GPM is different from the optimal $m$-GPM. (ii) if $m$ is the optimal piece number, then there is no additional piece can lower the error, making the optimal ($m+1$)-GPM is the same as the optimal $m$-GPM.  Now, no additional piece can be captured, hence $m$ is the optimal piece number. Appendix A.2 provides the full proof. ∎
Lemma 3.3 requires checking the results for every $\varepsilon$ and $x$, which is challenging as $\varepsilon\in[0,\infty)$ and $x\in\mathcal{D}$ are infinite sets. In practice, we restrict $\varepsilon$ within its generally meaningful domain, e.g. $\varepsilon\in[0,10)$, and employ Monte Carlo random sampling to generate a set of random pairs $(\mathcal{E,X})=\{(\varepsilon_{1},x_{1}),(\varepsilon_{2},x_{2}),\dots,(\varepsilon_{n},x_{n})\}$. If Lemma 3.3 holds for the $n$-size set $(\mathcal{E,X})$, the optimality of $m$-GPM is guaranteed with probability $1$ as $n\to\infty$ (Rubinstein, 1981).
(a) $m=4$.
(b) $m=5$.
Figure 2. Optimal $4$-GPM and $5$-GPM when $\varepsilon=1$ and $x=0.3$. They are identical as $m=3$ after merging redundant pieces.
Results for Monte Carlo sampling. Fixing $\mathcal{D}=\tilde{\mathcal{D}}=[0,1)$, for $\mathcal{L}=|y-x|$ and $\mathcal{L}=(y-x)^{2}$, the optimal $4$-GPM are identical as the optimal $3$-GPM for random $(\mathcal{E,X})$ with at least $n=10^{4}$. These optimal results align with TPM, but the optimal $p,l$ and $r$ values are different from existing instantiations.
Given the continuity of the objective functions w.r.t $\varepsilon$ and $x$, we posit that it is, in fact, the exact optimal. Therefore, when we say the optimal piece number is $m=3$, it implies the statistical optimality guaranteed by the Monte Carlo asymptotic technique with a strength of at least $n=10^{4}$ random samples.
Example 3.4.
For $[0,1)\to[0,1)$ and $\mathcal{L}=|y-x|$, Figure 2 shows two examples of the optimal $4$-GPM and optimal $5$-GPM when $\varepsilon=1$ and $x=0.3$. After merging redundant pieces, i.e. connected pieces with the same probability, they are the same as the $3$-GPM and fall into the TPM category.
3.2.3. Solution 3: Closed-form Instantiation
After determining the optimal $m$, we can derive the closed-form instantiation. If the optimal results exhibit $m=3$ and coincide with TPM, it facilitates analytical deduction for the closed-form optimal $p$, $l$ and $r$. Otherwise, numerical regression can be employed to obtain the closed-form instantiation. Figure 3 illustrates the workflow.
Analytical deduction. For TPM, the optimization procedure for solving $p_{i}$, $l_{i}$ and $r_{i}$ can be conducted analytically. The key observation is that there are only three variables: $p,l$, and $r$ in TPM. Due to the normalization constraint of probability, the central interval length $r-l$ can be replaced by its probability $p$. This reduces the solving for the optimal $p$ to a univariate optimization problem w.r.t. $p$, which can be solved by analyzing the first-order derivative. With the solved $p$, solving $l$ and $r$ also becomes a univariate optimization problem. Appendix B.4 provides the formalized process.
Numerical regression. For any $m$-GPM, we can obtain the closed-form $p_{i}$, $l_{i}$, and $r_{i}$ through numerical regression on their solved optimal values.
Assume that we want to find the closed-form optimal $p_{i}$. Given $(\mathcal{E},\mathcal{P}_{i})=\{(\varepsilon_{1},p_{i,1}),\dots,(\varepsilon_{n},p_{i,n})\}$, which contains the solved optimal $p_{i}$ for random $\varepsilon$, we aim to find the relationship between $\varepsilon$ and $p_{i}$. This relationship can be approximated by $\hat{p}_{i}=f(\varepsilon,\beta)$, where $f$ is a designed feature with $\beta$ as regression parameters.
Ideally, the designed feature $f$ matches the truth form of $p_{i}$. In this case, the regression result of $f$ on $(\mathcal{E},\mathcal{P}_{i})$ converges to the optimal closed-form $p_{i}$. If not, the regression result may not converge to the optimal. In practice, we can use heuristic forms of $f$ for tighter approximation. Due to the structure of the LDP constraint, we suggest choosing $p_{i}$ with a form of $\exp(\beta_{1}\varepsilon)$, allowing us to design $f=\exp(\beta_{1}\varepsilon)+\beta_{2}$.
Example 3.5.
For $\mathcal{L}=|y-x|$, $\mathcal{D}=\tilde{\mathcal{D}}=[0,1)$ and $m=3$, assume the probability $p$ has the form $\hat{p}=\exp(\beta_{1}\varepsilon)+\beta_{2}$. We use the scipy.curve_fit package to regress $p$ on $50$ random $\varepsilon$; then its regression result is $\hat{p}=\exp(\varepsilon/2)-0.07$ with a maximal error $\leq 10^{-2}$. Note that this result almost coincides with the ground-truth $p$ in Theorem 3.7.
Figure 3. Solving flow for the optimal closed form.
3.3. Closed Form for Classical Domain
The above framework for deriving the closed-form optimal GPM is applicable to any $\mathcal{D}\to\tilde{\mathcal{D}}$. Using this framework, this subsection provides instantiations for a common case: $\mathcal{D}=\tilde{\mathcal{D}}$.
Restricting domain. In real-world applications, the output domain is often required to match the input domain, i.e. $\mathcal{D}=\tilde{\mathcal{D}}$.666 $\tilde{\mathcal{D}}$ generally should be larger than or equal to $\mathcal{D}$ to ensure the mechanism’s meaningfulness; otherwise, it means some ranges in $\mathcal{D}$ will disappear after applying the mechanism. Furthermore, a concrete domain (e.g. $\mathcal{D}=[0,1)$) does not limit the generality, as $\mathcal{D}$ can be transformed to other domains through scaling and shifting operations. The theorem below characterizes the privacy and utility invariants under such transformations.
Theorem 3.6 (Transformation invariants).
Given a GPM $\mathcal{M}:\mathcal{D}\to\tilde{\mathcal{D}}$ satisfying $\varepsilon$-LDP, if domain $\mathcal{D}^{\prime}=c\mathcal{D}+d$ and $\tilde{\mathcal{D}}^{\prime}=c\tilde{\mathcal{D}}+d$ with $c>0$, then transformation
$\mathcal{T}(\mathcal{M}):x^{\prime}=cx+d,\ p^{\prime}_{i}=p_{i}/c,\ [l^{\prime}_{i},r^{\prime}_{i})=c[l_{i},r_{i})+d$
results in GPM $\mathcal{M}^{\prime}=\mathcal{T}(\mathcal{M}):\mathcal{D}^{\prime}\to\tilde{\mathcal{D}}^{\prime}$ having the same privacy level $\varepsilon$. (privacy invariant)
Meanwhile, if there is another GPM $\mathcal{M}_{\text{bad}}:\mathcal{D}\to\tilde{\mathcal{D}}$ with error $Err(x,\mathcal{M})\leq Err(x,\mathcal{M}_{\text{bad}})$ for a given $x$, then
$Err(x^{\prime},\mathcal{M}^{\prime})\leq Err(x^{\prime},\mathcal{M}^{\prime}_{\text{bad}}),$
i.e. $\mathcal{T}$ maintains the data utility ordering. (utility invariant)
Proof.
(Sketch) The privacy invariant is the same as applying a linear post-processing of DP to GPM. The utility invariant is due to $\mathcal{T}$ as a linear function on $\mathcal{D}$. Appendix A.3 provides the full proof of these two invariants. ∎
Theorem 3.6 allows us to discuss the optimality on a fixed input domain. If a mechanism’s input domain differs from $\mathcal{D}$, we can transform it to $\mathcal{D}$, and this transformation maintains the optimality.
Hypothesis 3.1.
For any domain $\mathcal{D}\to\mathcal{D}$, under absolute error and square error metrics, the optimal piecewise-based mechanism falls into $3$-GPM.
Support. We validated this hypothesis for $\mathcal{D}=[0,1)$ by performing Monte Carlo sampling on $10^{4}$ random $(\varepsilon,x)$ pairs (detailed in Section 3.2.2). Theorem 3.6 then extends this optimality to any $\mathcal{D}$. Given the continuity of the objective functions w.r.t. $\varepsilon$ and $x$, we posit that $m=3$ is indeed the exact optimal. To further support this hypothesis, Appendix B.3 outlines two directions for analytical proof and highlights the associated challenges.
Theorem 3.7.
If hypothesis 3.1 holds, then GPM $\mathcal{M}:[0,1)\to[0,1)$ with the following closed-form instantiation
$pdf[\mathcal{M}(x)=y]=\begin{dcases}p_{\varepsilon}&\text{if}\ y\in[l_{x,\varepsilon},r_{x,\varepsilon}),\cr p_{\varepsilon}/\exp{(\varepsilon)}&y\in[0,1)\,\backslash\,[l_{x,\varepsilon},r_{x,\varepsilon}),\end{dcases}$
where $p_{\varepsilon}=\exp(\varepsilon/2)$,
$[l_{x,\varepsilon},r_{x,\varepsilon})$ =$=\begin{dcases}[0,2C)&\text{if}\ x\in[0,C),\cr x+[-C,C)&\text{if }x\in[C,1-C),\cr[1-2C,1)&\text{otherwise},\cr\end{dcases}$
with $C=(\exp(\varepsilon/2)-1)/(2\exp(\varepsilon)-2)$, is optimal for $[0,1)\to[0,1)$ under the absolute error and square error metric.
Proof.
Provided by analytical deduction with $[a,b)=[\tilde{a},\tilde{b})=[0,1)$, $\mathcal{L}=|y-x|$ and $\mathcal{L}=|y-x|^{2}$ respectively. Appendix A.4 provides the full proof. ∎
This optimality on $[0,1)$ can be transformed to $\mathcal{M}^{\prime}:[a,b)\to[a,b)$ by applying $\mathcal{T}:x^{\prime}=(b-a)x+a,\,p^{\prime}_{\varepsilon}=p_{\varepsilon}/(b-a),\,[l^{\prime},r^{\prime})=(b-a)\cdot[l,r)+a$, while maintaining the optimality.
Example 3.8.
Figure 4 shows two examples of Theorem 3.7. When $x=0$ in the left figure, the optimal GPM has $p\approx 1.64$ and $[l,r)\approx[0,0.38)$. When $x=0.5$ in the right figure, the optimal GPM has $p\approx 1.64$ and $[l,r)\approx[0.31,0.69)$.
(a) $x=0$.
(b) $x=0.5$.
Figure 4. Optimal GPM (Theorem 3.7) when $\varepsilon=1$, $x=0$ and $x=0.5$.
MSE analysis. We can calculate the mean squared error (MSE) of the optimal GPM in Theorem 3.7 as
[…]
which leads to the results in Appendix B.5.
The theoretical MSE allows us to analytically compare the optimal GPM with existing mechanisms. For mechanisms defined on $\mathcal{D}=[0,1)$, e.g. SW (Li et al., 2020), we can compare their MSE with the above result. For mechanisms defined on other domains, e.g. PM (Wang et al., 2019) on $\mathcal{D}=[-1,1)$, we can transform the optimal GPM to $[-1,1)$ and compare their MSE. Detailed comparison with them is presented in the evaluation section.
4. Optimal GPM for Circular Domain
This section presents the optimal GPM for the circular domain, another type of bounded domain. Circular domains are widely used in cyclic data such as time, angle, and compass direction. However, none of the existing piecewise-based mechanisms consider this type of domain, limiting their applicability.
Different meanings of distance. In the circular domain $[0,2\pi)$, the distance between two elements differs from that in the classical domain $[0,2\pi)$. For example, if we denote the distance between $x$ and $y$ in the circular domain as $\mathcal{L}_{\text{mod}}(y,x)=|y-x|$, then it implies $\mathcal{L}_{\text{mod}}(2\pi,0)=0$ and $\mathcal{L}_{\text{mod}}(3\pi/2,0)=\pi/2$, which are different from $\mathcal{L}(y,x)=|y-x|$ in the classical domain.  The biggest difference is that there are no endpoints in the circular domain. This unique property makes the mechanisms designed for the classical domain not directly suitable for the circular domain. Although we can “flatten” the circular domain to the classical domain, this conversion changes the distance between elements, leading to data utility loss.
Formally, in the circular domain $[0,2\pi)$, the distance metric $\mathcal{L}_{\text{mod}}(y,x)$ has the following relationship with $\mathcal{L}(y,x)$ in the classical domain:
$\mathcal{L}_{\text{mod}}(y,x)=\min\big{(}\mathcal{L}(y,x),\mathcal{L}(y,2\pi-x)\big{)},$
i.e. the distance between $y$ and $x$ is the smaller one between two arcs from $y$ to $x$. Under this distance metric, finding the optimal $m$-GPM for the circular domain is to solve $\mathcal{M}:[0,2\pi)\to[0,2\pi)$ such that
[…]
and use the solved optimal $x$-independent $p_{i}$ to determine $l_{i,x},r_{i,x}$ for any given $x$:
[…]
We show that these two problems can be reduced to those in the classical domain, thereby enabling the usage of existing results.
4.1. Reduced Forms
Similar to the classical domain, the min-max objective of Formulation (4) also has a closed-form solution, which reduces the problem to the classical domain.
Lemma 4.1.
The objective of Formulation (4) can be reduced to
$\min_{p_{i},l_{i,x},r_{i,x}}\int_{0}^{2\pi}\mathcal{L}(y,\pi)\mathcal{P}_{\mathcal{M}(\pi)}\mathrm{d}y.\$
Proof.
(Sketch) We prove it by showing that, for any fixed $y$, $\max_{x}\mathcal{L}_{\text{mod}}(y,x)=\mathcal{L}_{\text{mod}}(y,\pi)=\mathcal{L}(y,\pi)$, i.e. the maximum distance between $y$ and $x$ is achieved at a unique $x=\pi$ for any $y$. Appendix A.5 provides the full proof. ∎
Following this reduction, the optimal results in the classical domain $[0,2\pi)$ can be applied to determine the optimal $p_{i}$.
For Formulation (5) to solve $l_{i,x}$ and $r_{i,x}$, we shift the domain $[0,2\pi)$ by $\pi-x$. This is a trick operation as $\mathcal{L}_{\text{mod}}(y,x)$ is transformed to the following form:
$\begin{split}&\mathcal{L}_{\text{mod}}(y+\pi-x,x+\pi-x)=\mathcal{L}_{\text{mod}}(y+\pi-x,\pi)\\ =&\min\left(\mathcal{L}(y+\pi-x,\pi),\mathcal{L}(y+\pi-x,2\pi-\pi)\right)\\ =&\mathcal{L}(y+\pi-x,\pi),\end{split}$
which transforms $\mathcal{L}_{\text{mod}}$ to $\mathcal{L}$ at $x=\pi$. Then, the optimization problem in the shifted domain becomes
$\min_{l_{i,x},r_{i,x}}\int_{\pi-x}^{3\pi-x}\mathcal{L}(y+\pi-x,\pi)\mathcal{P}_{\mathcal{M}(\pi)}\mathrm{d}y.$
It is a problem in the classical domain $[\pi-x,3\pi-x)$. We can obtain the closed-form optimal $l_{i,x}$ and $r_{i,x}$ by applying the results of the classical domain. Since the obtained $l_{i,x}$ and $r_{i,x}$ depends on $x$ in the shifted domain, we shift them back to the circular domain using
$l_{i,x}^{\text{mod}}$ $=l_{i,x}-(\pi-x)\mod 2\pi,$
$r_{i,x}^{\text{mod}}$ $=r_{i,x}-(\pi-x)\mod 2\pi.$
Transformation invariants ensure their optimality in the circular domain. Figure 5 summarizes the above two reductions to solve the optimal $p_{i}$, $l_{i,x}^{\text{mod}}$ and $r_{i,x}^{\text{mod}}$ in the circular domain.
Figure 5. Reduced forms of solving the optimal $p_{i}$, $l_{i,x}^{\text{mod}}$ and $r_{i,x}^{\text{mod}}$. Optimizations under $\mathcal{L}_{\text{mod}}$ can be reduced to those under $\mathcal{L}$.
4.2. Closed Form for Circular Domain
By applying the above reductions and following the same steps as in the classical domain, we can obtain the optimal GPM for the circular domain.
Theorem 4.2.
If Hypothesis 3.1 holds, then GPM $\mathcal{M}:[0,2\pi)\to[0,2\pi)$ with the following closed-form instantiation
$pdf[\mathcal{M}(x)=y]=\begin{dcases}p_{\varepsilon}&\text{if}\ y\in[l_{x,\varepsilon}^{\textup{mod}},r_{x,\varepsilon}^{\textup{mod}}),\cr p_{\varepsilon}/\exp{(\varepsilon)}&y\in[0,2\pi)\,\backslash\,[l_{x,\varepsilon}^{\textup{mod}},r_{x,\varepsilon}^{\textup{mod}}),\end{dcases}$
where $p_{\varepsilon}=\frac{1}{2\pi}\exp(\varepsilon/2)$,
$l_{x,\varepsilon}^{\textup{mod}}$ $=\left(x-\pi\frac{\exp(\varepsilon/2)-1}{\exp(\varepsilon)-1}\right)\mod 2\pi,$
$r_{x,\varepsilon}^{\textup{mod}}$ $=\left(x+\pi\frac{\exp(\varepsilon/2)-1}{\exp(\varepsilon)-1}\right)\mod 2\pi,$
is optimal for the circular domain under the absolute error and square error metric.
Proof.
Combination of the reduced forms, the results of the classical domain, Lemma 4.1, and transformation invariants leads to the conclusion. ∎
Compared to the optimal GPM for the classical domain in Theorem 3.7, the key difference is that the mechanism in the circular domain allows $[l_{x,\varepsilon}^{\textup{mod}},r_{x,\varepsilon}^{\textup{mod}})$ to span the 00 or $2\pi$ boundary, significantly reducing the error. We also observe that their instantiations of $p_{\varepsilon},l_{x,\varepsilon},r_{x,\varepsilon}$ are connected through transformation invariants. For instance, moving from the classical domain $[0,1)$ to the circular domain $[0,2\pi)$, $p_{\varepsilon}=\exp(\varepsilon/2)$ transforms to $p_{\varepsilon}=\frac{1}{2\pi}\exp(\varepsilon/2)$, reflecting the ratio of $[0,1)$ to $[0,2\pi)$.
Example 4.3.
Figure 6 shows two examples of Theorem 4.2 when $\varepsilon=1$. For $x=0$ in the left figure, it samples the output $y$ from $[1.62\pi,2\pi)\cup[0,0.38\pi)$ with probability density $0.26$ and from $[0.38\pi,1.62\pi)$ with probability density $0.09$.
(a) $x=0$.
(b) $x=\pi/4$.
Figure 6. Optimal GPM (Theorem 4.2) for the circular domain with $\varepsilon=1$. Here $[0,2\pi)$ is wrapped into a circle. Angle values in the red arc (centered at $x$) have a higher sampling pdf.
MSE analysis. The MSE of the optimal GPM in Theorem 4.2 needs a separate analysis due to the circular domain. The biggest difference from the classical domain is that there exist no endpoints, i.e. fixed farthest points, in the circular domain. Without loss of generality, assume :-$\mathcal{L}\coloneq|y-x|^{2}$ and $x>\pi$, the farthest distance from $x$ is always $\pi$, i.e. from $x$ to $x-\pi$. If we shift the data domain by $\pi-x$, then point $x-\pi$ is mapped to 00. This domain shift does not change the value of $\mathcal{L}$, but $x$ now locates at $\pi$.  Therefore, the MSE of the optimal GPM in Theorem 4.2 has an identical value for all $x$ in the circular domain, which is
$\begin{split}&MSE(\mathcal{M}(x),x)=MSE(\mathcal{M}(\pi),\pi)\\ =&2\left(\int_{0}^{l_{\pi,\varepsilon}^{\textup{mod}}}(y-\pi)^{2}\frac{p_{\varepsilon}}{\exp(\varepsilon)}\mathrm{d}y+\int_{l_{\pi,\varepsilon}^{\textup{mod}}}^{\pi}(y-\pi)^{2}p_{\varepsilon}\mathrm{d}y\right).\end{split}$
Calculating the above integral, we can obtain the MSE of the optimal GPM in the circular domain.
Theorem 4.4.
The optimal GPM in Theorem 4.2 has an identical MSE for all $x$ in the circular domain, which is
$MSE(\mathcal{M}(x),x)=\frac{2}{3}\left((\pi^{3}-C^{3})\frac{p_{\varepsilon}}{\exp(\varepsilon)}+C^{3}p_{\varepsilon}\right),$
where $C=\pi(\exp(\varepsilon/2)-1)/(\exp(\varepsilon)-1)$.
Proof.
We have shown that the MSE at $x$ is the same as that at $\pi$. Then the calculation is straightforward by plugging in the values of $p_{\varepsilon}$ and $l_{\pi,\varepsilon}^{\textup{mod}}$. ∎
If we do not consider $\mathcal{L}_{\text{mod}}$ and directly apply the optimal GPM in the classical domain (or SW and PM mechanisms) to the circular domain, i.e. flatten the circular domain to the classical domain $[0,2\pi)$ and consider $\mathcal{L}$, the MSE will varies for different $x$. In the flattened domain, the worst-case MSE is at $x=0$ or $x=2\pi$ and the best-case MSE is at $x=\pi$. Therefore, the optimal GPM in Theorem 4.2 always has a lower MSE than “flattened” mechanisms. The evaluation section will further demonstrate this.
5. Distribution and Mean Estimation
This section applies our mechanisms to support the commonly used distribution and mean estimation. We will show that our mechanisms also provide optimality for these estimation tasks.
Assume a set of users with sensitive data $\mathcal{X}=\{x_{1},x_{2},\dots,x_{n}\}$. They apply $\mathcal{M}$ to produce randomized outputs $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$. The data collector then estimates the distribution and mean of $\mathcal{X}$ using $\mathcal{Y}$. Specifically, the collector uses the values in $\mathcal{Y}$ as it is, i.e. without knowing and applying any post-processing based on prior knowledge of $\mathcal{X}$.
5.1. Distribution Estimation
To estimate the distribution of $\mathcal{X}$ from $\mathcal{Y}$, the collector must discretize the continuous domain $\mathcal{D}$ into $k$ bins $B_{1},B_{2},\dots,B_{k}$. Each bin $B_{j}$’s probability is then estimated by the proportion of $y_{i}$ that falls within $B_{j}$. Probabilities of all bins together form the estimated distribution ^$\hat{\mathcal{F}}_{B}$ using $\mathcal{Y}$. This estimation’s accuracy is measured by the distance between the estimated distribution ^$\hat{\mathcal{F}}_{B}$ and the true distribution $\mathcal{F}_{B}$ from $\mathcal{X}$ (Li et al., 2020).
Note that the bin size can impact the estimation accuracy due to the rounding of the bins. However, it is actually a hyperparameter that does not inherently affect the estimation accuracy if we have a sufficiently large number of bins, as the support of ^$\hat{\mathcal{F}}_{B}$ (i.e. non-zero bins) converges to $\mathcal{Y}$ and the support of $\mathcal{F}_{B}$ converges to $\mathcal{X}$, i.e.
$\lim_{k\to\infty}\mathrm{supp}(\hat{\mathcal{F}}_{B})=\mathcal{Y},\quad\lim_{k\to\infty}\mathrm{supp}(\mathcal{F}_{B})=\mathcal{X}.$
Our mechanisms guarantee the optimal error between $\mathcal{X}$ and $\mathcal{Y}$ for each $x_{i}$. This ensures their optimality among GPM when applied to distribution estimation.
Some other statistical estimations are special applications of distribution estimation, such as range and quantiles (Li et al., 2020) to estimate a specific part of the distribution. Our mechanisms have optimality for these estimations as well.
5.2. Mean Estimation
To estimate the mean of $\mathcal{X}$ from $\mathcal{Y}$, the collector uses the estimator $\hat{\mu}=\sum_{i=1}^{n}y_{i}/n$. The accuracy of this estimator is measured by $|\hat{\mu}-\mu|$, where $\mu$ is the true mean of $\mathcal{X}$. Our mechanisms guarantee the optimal error between each $x_{i}$ and $y_{i}$, which tends to give the smallest $|\hat{\mu}-\mu|$ for any $n$ among GPM.
Typically, a mean estimator may also need to be unbiased, i.e. $\mathrm{E}[\hat{\mu}]=\mu$. This constraint translates to $\mathrm{E}[y_{i}]=\mathrm{E}[\mathcal{M}(x_{i})]=x_{i}$.777 Actually, unbiasedness for numerical data is not as important as for categorical data, as it is for a single data point $x_{i}$. For numerical data with infinite possible data points, the probability of having a specific $x_{i}$ is zero. This is unachievable for same-domain mapping $\mathcal{M}:\mathcal{D}\to\mathcal{D}$ on classical domains, as the endpoints of $\mathcal{D}$ cannot be the mean value (or center) of any distribution over $\mathcal{D}$. For example, for any LDP mechanism $\mathcal{M}:[0,1)\to[0,1)$, distribution of $\mathcal{M}(0)$ can not be unbiased. So the mechanism in Theorem 3.7 is biased.
Unbiased mean estimation. Note that an unbiased mean estimator can be achieved by enlarging the output domain $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$. Mathematically, this involves incorporating the unbiasedness constraint $\mathrm{E}[\mathcal{M}(x)]=x$ into optimization problems for solving $\mathcal{M}$. Following the same optimization process as in the classical domain, we hypothesize that the $3$-GPM remains optimal for domain $\tilde{\mathcal{D}}_{\varepsilon}$.
Hypothesis 5.1.
For any domain $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$, where $\tilde{\mathcal{D}}_{\varepsilon}$ is a variable w.r.t $\varepsilon$, and under absolute error and square error metrics, the optimal piecewise-based mechanism falls into $3$-GPM.
Under the $3$-GPM, an unbiased mechanism $\mathcal{M}$ with a variable output domain $\tilde{\mathcal{D}}_{\varepsilon}$ can be analytically derived by incorporating the unbiasedness constraint. As a complement to Theorem 3.7, we provide Theorem 5.1 for mean estimation [^8] However, for GPM, unbiasedness does not always result in better data utility for mean estimation, as it comes at the cost of a larger output domain, which leads to higher errors (as shown in Figure 17). in the classical domain.
Theorem 5.1.
Denote $\tilde{\mathcal{D}}_{\varepsilon}=[-C,C+1)$ with $C=(\exp(\varepsilon/2)+1)/(\exp(\varepsilon/2)-1)$. If Hypothesis 5.1 holds, then among the unbiased GPM $\mathcal{M}:\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$ (i.e. $\mathrm{E}[\mathcal{M}(x)]=x$), closed form
$pdf[\mathcal{M}(x)=y]=\begin{dcases}p_{\varepsilon}&\text{if}\ y\in[l_{x,\varepsilon},r_{x,\varepsilon}),\cr p_{\varepsilon}/\exp{(\varepsilon)}&y\in\tilde{\mathcal{D}}_{\varepsilon}\,\backslash\,[l_{x,\varepsilon},r_{x,\varepsilon}),\end{dcases}$
where $p=\exp(\varepsilon/2)/(2C+1)$,
$l_{x,\varepsilon}$ $=\frac{C+1}{2}\cdot x-\frac{(3C+1)(C-1)}{4C},$
$r_{x,\varepsilon}$ $=\frac{C+1}{2}\cdot x+\frac{(C+1)(C-1)}{4C}.$
is optimal for $[0,1)\to\tilde{\mathcal{D}}_{\varepsilon}$ and the square error metric.
Proof.
Instantiations of $C$, $p$, $l_{x,\varepsilon}$, and $r_{x,\varepsilon}$ are derived by analytical deduction. Appendix A.6 proves the unbiasedness. ∎
In the context of the circular domain, the $\mathcal{M}$ in Theorem 4.2 is unbiased, as the distribution of $\mathcal{M}$ is always centered at $x$. This property is also illustrated in Figure 6.
Table 2 summarizes the optimal GPM for both distribution and mean estimation in three domain types. Theorems use a concrete domain $\mathcal{D}$ for the sake of clarity. We do not give closed-form optimal GPM for distribution estimation on $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$ because $\tilde{\mathcal{D}}_{\varepsilon}$ can not be restricted by constraints as in mean estimation.
Table 2. Optimal distribution and mean estimation in three domain types under GPM.
Distribution Mean
Classical ($\mathcal{D}\to\mathcal{D}$) Theorem 3.7 Theorem 3.7 (biased)
Circular domain Theorem 4.2 Theorem 4.2 (unbiased)
Classical ($\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$) Theorem 5.1 (unbiased)
6. Discussion and Extension
Minimize Error at a Specific $x$
Our proposed optimal GPMs are designed to minimize the worst-case error over the whole domain. However, the framework can be used to minimize the error at any specific $x$. This is useful when the data distribution is concentrated around a specific data point.
Formally, assume the data distribution is concentrated around $x_{0}$, and we want to minimize the error at $x_{0}$. The optimization problem in Lemma 3.2 for solving $p_{i}$ can be modified to
$\min_{p_{i},l_{i,x},r_{i,x}}\int_{\tilde{\mathcal{D}}}\mathcal{L}(y,x_{0})\mathcal{P}_{\mathcal{M}(x_{0})}\mathrm{d}y.$
This optimization problem generally leads to a different $p_{i}$ from the optimal GPM for the worst-case error.
Example 6.1.
Assume $\mathcal{D}\to\tilde{\mathcal{D}}=[0,1)\to[0,1)$, :-$\mathcal{L}\coloneq|y-x|$, and the data distribution is concentrated around $x_{0}=0.2$. Solving the above optimization problem with $\varepsilon=1$ gives probability (of the second piece) $p_{2}=1.54$ and $Err(x_{0})=0.241$. In contrast, the optimal GPM for the worst-case error uses $p_{2}=1.64$, as shown in Figure 4, which gives $Err(x_{0})=0.243$.
Importantly, optimizing for a specific $x_{0}$ does not leak information about $x_{0}$, as the mechanism (i.e. $p_{i}$, $l_{i,x}$, and $r_{i,x}$) still contains no information about $x_{0}$. Moreover, observing $Err(x)$ at all $x$ does not reveal $x_{0}$, as the error at $x_{0}$ is not necessarily the smallest. Therefore, the adversary cannot infer $x_{0}$ from observing the mechanism.
An Extension: 2D Polar Coordinates
Polar coordinates are widely used in relative location representation, e.g. navigation systems that have the locations of surrounding objects relative to it. Our proposed optimal GPM for the classical and circular domain can be combined and extended to polar coordinates for collecting such data under LDP.
Privacy. A polar coordinate data is represented by a 2D tuple $(x_{1},x_{2})\in[0,d)\times[0,2\pi)$, where $x_{1}$ is the distance from the pole and $x_{2}$ is the angle from the polar axis. The first dimension is linear, while the second is naturally circular, thus we can combine the optimal GPM for both domains to provide LDP for such data.
Utility. Our mechanisms guarantee the optimal error for each dimension. Therefore, if we use $\mathcal{L}_{\text{2D}}\coloneq\mathcal{L}(y_{1},x_{1})+\mathcal{L}_{\text{mod}}(y_{2},x_{2})$ as the error metric for the polar coordinate data and optimally assign the privacy parameter $\varepsilon$ to each dimension, the optimal GPM preserves the optimal error.
Optimal assignment of $\varepsilon$. Formally, we want to assign $\varepsilon=\varepsilon_{1}+\varepsilon_{2}$ to $x_{1}$ and $x_{2}$ respectively to minimize the worst-case error in 2D polar coordinates. This error is the sum of the worst-case error for $x_{1}$ and $x_{2}$ under $\mathcal{L}$ and $\mathcal{L}_{\text{mod}}$ respectively. Therefore, the optimal assignment of $\varepsilon$ can be derived by solving
$\min_{\varepsilon_{1},\varepsilon_{2}}Err_{\text{1,wor}}(\varepsilon_{1})+Err_{\text{2,wor}}(\varepsilon_{2}),$
where $Err_{\text{1,wor}}(\varepsilon_{1})$ and $Err_{\text{2,wor}}(\varepsilon_{2})$ are the worst-case error for the classical domain $[0,d)$ and the circular domain $[0,2\pi)$ respectively. Closed-form $Err_{\text{1,wor}}(\varepsilon_{1})$ and $Err_{\text{2,wor}}(\varepsilon_{2})$ can be derived from instantiations of the mechanism and the error metric. Then the above optimization problem gives the optimal $\varepsilon_{1}$ and $\varepsilon_{2}$. Appendix B.6 provides details for solving this optimization problem.
Example 6.2.
Figure 7 shows two examples of the optimal GPM for 2D polar coordinates in $[0,1)\times[0,2\pi)$ with $\varepsilon=1+2\pi$. The green point represents the sensitive data, the optimal GPM samples the output from the pink area with a higher probability.
(a) $(x_{1},x_{2})=(0.5,2\pi)$.
(b) $(x_{1},x_{2})=(1,\pi/2)$.
Figure 7. Optimal GPM for 2D polar coordinates when $\varepsilon=1+2\pi$ and $\mathcal{L}=|y-x|^{2}$ under $\mathcal{L}_{\text{2D}}\coloneq\mathcal{L}(y_{1},x_{1})+\mathcal{L}_{\text{mod}}(y_{2},x_{2})$.
7. Evaluations
(a) Privacy parameter $\varepsilon=2$.
(b) Privacy parameter $\varepsilon=4$.
Figure 8. Whole-domain error comparison in the classical domain with error metric $\mathcal{L}=|y-x|$.
(c) Distance metric $\mathcal{L}=|y-x|$.
(d) Distance metric $\mathcal{L}=|y-x|^{2}$.
Figure 9. Worst-case error comparison in the classical domain.
(a) Privacy parameter $\varepsilon=2$.
(b) Privacy parameter $\varepsilon=4$.
Figure 10. Whole-domain error comparison in the circular domain with error metric $\mathcal{L}=|y-x|^{2}$.
(c) Distance metric $\mathcal{L}=|y-x|$.
(d) Distance metric $\mathcal{L}=|y-x|^{2}$.
Figure 11. Worst-case error comparison in the circular domain.
This section evaluates the theoretical and experimental data utility of our methods by comparing them with existing instantiations and their variants:
  • OGPM: closed-form optimal GPM (Theorem 3.7 and 4.2 for the classical and circular domain, respectively).
  • OGPM-U: unbiased closed-form optimal GPM (Theorem 5.1) for mean estimation in the classical domain.
  • PM (Wang et al., 2019), SW (Li et al., 2020), and their post-processed versions: PM is the first TPM designed for mean estimation, while SW is designed for distribution estimation. Both mechanisms output enlarged domains but can be post-processed by truncating outputs to the input domain. These post-processed versions are referred to as T-PM and T-SW for convenience.
  • PM-C and SW-C: the compressed versions of PM and SW for $\mathcal{D}\to\mathcal{D}$. For the best potential of PM and SW, we adapt them to $\mathcal{D}\to\mathcal{D}$ as PM-C and SW-C by linearly compressing their output domain $\tilde{\mathcal{D}}_{\varepsilon}$ to $\mathcal{D}$, i.e. transformation invariants, which maintains the privacy level.
We also compare OGPM’s expected error with non-piecewise-based mechanisms that can be applied to bounded domains:
  • Variants of the Laplace mechanism: including the staircase mechanism (Geng and Viswanath, 2014), Laplace mechanism with post-processing by truncation (T-Laplace), and the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020), which designs a bounded Laplace-shape distribution.
  • Purkayastha mechanism (Weggenmann and Kerschbaum, 2021): a mechanism for directional data on spheres $\mathbb{S}^{n-1}$. When $n=2$, it is a counterpart of OGPM in the circular domain.
We omit comparison with PTT (Ma et al., 2024) because it does not provide a concrete method to find a closed-form mechanism. We use $\mathcal{D}=[0,1)$ as the classical domain; this does not change their data utility ordering by the transformation invariant.  PM and SW can only be applied to the classical domain, so when evaluating them in the circular domain, we “flatten” the circular domain to the classical domain $[0,2\pi)$ and apply them to the flattened domain.
The first subsection presents the comparison of expected errors, followed by the distribution and mean estimations on real-world datasets in the second subsection.
7.1. Expected Errors
GPM’s data utility under distance metric $\mathcal{L}$ is measured by the expected error $Err(x)$ in Formula (1):
$Err(x)=\int_{\tilde{\mathcal{D}}}\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}\mathrm{d}y,$
where $x$ is the sensitive input and $y$ is the output of $\mathcal{M}$. Based on $Err(x)$, two types of error need to be considered:
  • whole-domain error: $Err(x)$ values for the whole domain, i.e. for all $x\in\mathcal{D}$.
  • worst-case error: the largest $Err(x)$ value among the whole domain. PM (Wang et al., 2019) and PTT (Ma et al., 2024) use this error to evaluate data utility.
We have proved that the worst-case error of the classical domain is from the endpoints, so this error is actually the error at $x=0$ and $x=1$ for the classical domain. For the circular domain, the worst-case error is at $x=\pi$.
7.1.1. Comparison with PM-C and SW-C
PM-C and SW-C exhibit the best potential of PM and SW, so we compare OGPM with them in the first place. The comparisons are conducted under the classical domain and the circular domain, respectively.
Classical Domain. Figure 9 shows the comparison of the whole-domain error in the classical domain. We use distance metric $\mathcal{L}=|y-x|$ and set $\varepsilon=2$ and $\varepsilon=4$ for the comparison. It can be seen that OGPM consistently has the lowest error across all $x$ values.  For all mechanisms, the expected error achieves the maximal value at the endpoints and the minimal value at the midpoint. At small $\varepsilon$ values, their errors are similar due to the strong randomness (privacy constraint). At larger $\varepsilon$ values, OGPM’s error has a significant advantage over PM-C and SW-C, especially between the endpoints and the midpoint.  Statistically, under $\mathcal{L}=|y-x|$ with $\varepsilon=2$, OGPM’s average error is $94.2\%$ of PM-C and $92.3\%$ of SW-C across the whole domain. When $\varepsilon=4$, OGPM’s average error is $90.5\%$ of PM-C and $74.7\%$ of SW-C. More comparisons under smaller $\varepsilon$ values are provided in Appendix B.7.
Figure 9 shows the comparison of the worst-case error w.r.t $\varepsilon$ in the classical domain. The error is measured by two distance metrics: $\mathcal{L}=|y-x|$ and $\mathcal{L}=|y-x|^{2}$. It can be seen that the error of all mechanisms decreases with $\varepsilon$, but OGPM and PM-C decreases faster than SW-C.  For $\mathcal{L}=|y-x|$, OGPM has almost the same worst-case error as PM-C; for $\mathcal{L}=|y-x|^{2}$, OGPM’s error is slightly smaller than PM-C. For both metrics, OGPM’s worst-case error is the lowest across all $\varepsilon$ values.  Statistically, under $\mathcal{L}=|y-x|^{2}$, OGPM’s average error is $89.9\%$ of PM-C and $61.7\%$ of SW-C.
Circular Domain. Figure 11 shows the comparison of the whole-domain error in the circular domain. We use distance metric $\mathcal{L}=|y-x|^{2}$ and set $\varepsilon=2$ and $\varepsilon=4$ for the comparison. It can be seen that OGPM consistently has the lowest error across all $x$ values, and the error is stable across $x$, which is consistent with the theoretical analysis in Theorem 4.4.  For PM-C and SW-C, which treat the circular domain as the classical domain, their errors vary with $x$ and are higher than OGPM’s error, especially near the endpoints.  Statistically, under $\mathcal{L}=|y-x|^{2}$ with $\varepsilon=2$, OGPM’s average error is $47.5\%$ of PM-C and $43.0\%$ of SW-C across the whole domain. When $\varepsilon=4$, OGPM’s average error is $41.3\%$ of PM-C and $24.6\%$ of SW-C.
Figure 11 shows the comparison of the worst-case error w.r.t $\varepsilon$ in the circular domain. The error is measured by two distance metrics: $\mathcal{L}=|y-x|$ and $\mathcal{L}=|y-x|^{2}$. Similar to the classical domain, OGPM has the lowest worst-case error across all $\varepsilon$ values, and the advantage is more significant, especially for small $\varepsilon$ values.  Statistically, under $\mathcal{L}=|y-x|$, OGPM’s average error is $50.0\%$ of PM-C and $41.6\%$ of SW-C across the range of $\varepsilon$. Under $\mathcal{L}=|y-x|^{2}$, OGPM’s average error is $22.4\%$ of PM-C and $15.4\%$ of SW-C.
The above comparisons show that OGPM has the lowest expected error in both the classical and circular domains. The advantage of OGPM is more significant in the circular domain, where the error is stable across $x$.
7.1.2. Comparison with PM and SW
Figure 12 presents the comparison of the whole-domain error in the classical domain for the original PM and SW mechanisms, along with their post-processed versions, T-PM and T-SW. For a fair comparison, OGPM is adapted to the domain $\mathcal{D}=[-1,1)$ to match PM’s design, while SW and OGPM remain consistent with $\mathcal{D}=[0,1)$. The post-processing of PM and SW involves truncating their outputs in the enlarged domain to the input domain, i.e. applying $\mathcal{I}\circ\mathcal{M}(x)$, where $\mathcal{I}:\tilde{\mathcal{D}}\to\mathcal{D}$ is the truncation operator. We use the distance metric $\mathcal{L}=|y-x|^{2}$ and set $\varepsilon=2$ for the comparison among these five mechanisms. It can be observed that OGPM consistently achieves the lowest error across all $x$ values, with a more significant advantage compared to the comparison with PM-C and SW-C. This is because the original PM and SW output larger domains, resulting in higher errors. Meanwhile, T-PM reduces the error of PM more effectively than T-SW reduces the error of SW, as the original PM has a more enlarged output domain than SW, making truncation more impactful. This comparison highlights OGPM’s error advantage over the original PM, SW, and their post-processed versions when applied to their respective data domains.
(a) Comparison with PM.
(b) Comparison with SW.
Figure 12. Whole-domain error comparison with PM and SW on their data domains (i.e. $\mathcal{D}=[-1,1)$ and $\mathcal{D}=[0,1)$, respectively) when $\varepsilon=2$.
(a) Privacy parameter $\varepsilon=2$.
(b) Privacy parameter $\varepsilon=4$.
Figure 13. Whole-domain error comparison with the staircase mechanism (Geng and Viswanath, 2014), T-Laplace and B-Laplace mechanisms (Holohan et al., 2020) in the classical domain with error metric $\mathcal{L}=|y-x|$.
7.1.3. Comparison with the Staircase Mechanism, T-Laplace, and B-Laplace
In addition to piecewise-based mechanisms, the Laplace mechanism and its variants can also be applied to the classical domain to achieve LDP. Among these, the staircase mechanism (Geng and Viswanath, 2014) claims to be optimal under certain assumptions. For the input domain $\mathcal{D}=[0,1)$ (i.e. sensitivity $\Delta=1$) and error metric $\mathcal{L}=|y-x|$, its expected error is given by Theorem 3 in (Geng and Viswanath, 2014): $\exp(\varepsilon/2)/(\exp(\varepsilon)-1)$.  Another approach involves using the Laplace mechanism with truncation (Holohan et al., 2020), referred to here as T-Laplace for convenience. T-Laplace preserves the privacy guarantees of the Laplace mechanism while reducing the expected error, particularly for data points near the endpoints or for small $\varepsilon$ values. Additionally, the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020) introduces a redesigned bounded Laplace-shaped distribution tailored for bounded domains.999 Appendix B.8 provides details on the expected error of B-Laplace.
Figure 14. Worst-case error comparison (continued from Figure 13).
Figure 15. Comparison with the Purkayastha mechanism (Weggenmann and Kerschbaum, 2021) for sphere $\mathbb{S}^{n-1}$.
Figure 13 compares the whole-domain error in the classical domain $\mathcal{D}=[0,1)$ for the staircase mechanism, T-Laplace, and B-Laplace. These mechanisms exhibit distinct error patterns across the domain.  For the staircase mechanism, the error remains constant, as it is determined by a fixed staircase distribution and is independent of $x$. For T-Laplace, the error reaches its maximum at the midpoint and its minimum at the endpoints, as truncation favors the endpoints. For instance, when $x=0$, it is error-free with a probability of $1/2$, due to the symmetry of the Laplace distribution around 00. For B-Laplace, the error trend varies with $\varepsilon$. When $\varepsilon=2$, the error decreases with $x$ and reaches its minimum at the midpoint, whereas for $\varepsilon=4$, the error increases with $x$ and peaks at the midpoint. Despite these differing error patterns, OGPM generally achieves lower errors than the staircase mechanism, T-Laplace, and B-Laplace across the whole domain.
Figure 15 compares the worst-case error w.r.t. $\varepsilon$ in the classical domain. OGPM consistently achieves the lowest worst-case error across all $\varepsilon$ values. For small $\varepsilon$, T-Laplace and B-Laplace exhibit a significant advantage over the staircase mechanism; however, this advantage diminishes as $\varepsilon$ increases. At larger $\varepsilon$ values, the error of the staircase mechanism approaches that of OGPM.
7.1.4. Comparison with the Purkayastha Mechanism
The paper “Differential Privacy for Directional Data” (Weggenmann and Kerschbaum, 2021) introduces two mechanisms for data on spheres $\mathbb{S}^{n-1}$: the VMF mechanism (ensuring indistinguishability of any two points with distance through the sphere) and the Purkayastha mechanism (ensuring indistinguishability of any two points with distance along the sphere). When $n=2$, the sphere $\mathbb{S}^{1}$ corresponds to a circle, making the Purkayastha mechanism a counterpart of OGPM in the circular domain. Therefore, we compare them in the circular domain.101010 We omit the comparison with the VMF mechanism also because (i) it has been shown that the Purkayastha mechanism outperforms the VMF mechanism (with the same sensitivity $\Delta_{\measuredangle}=\pi$ for sphere $\mathbb{S}^{1}$, e.g. Figure 5 and 10 in (Weggenmann and Kerschbaum, 2021)), and (ii) the expected error of the VMF mechanism lacks a closed-form expression (Theorem 17 in (Weggenmann and Kerschbaum, 2021)), making it complex to compute.
Figure 15 presents the comparison of the expected error in the circular domain between OGPM and the Purkayastha mechanism. The expected error of the Purkayastha mechanism is derived using the closed-form expressions in Theorem 19 and 22 of (Weggenmann and Kerschbaum, 2021), with $\kappa=\varepsilon/\Delta_{\measuredangle}$. Since the errors of both mechanisms are $x$-independent in the circular domain, it suffices to compare their worst-case errors. The results demonstrate that OGPM consistently outperforms the Purkayastha mechanism, achieving significantly lower errors.
7.2. Distribution and Mean Estimations
This section compares the experimental data utility of the mechanisms in distribution and mean estimations.
7.2.1. Setup
We choose the MotionSense dataset (noa, 2018; Malekzadeh et al., 2019) for the evaluation. It contains smartphone sensor data recorded during various human activities. Specifically, we use the data from the first three files, encompassing a total of 6 15961596\,1596 159 data entries. We focus on two types of sensors:
  • Accelerometer (linear data): We normalize the dataset to $[0,1)$ for the classical domain.
  • Attitude sensor (angular data): We use this dataset for the circular domain.
Upon applying each LDP mechanism to these datasets, we evaluate the accuracy of distribution and mean estimations on them. For distribution estimation, we divide the domain into $k=50$ bins and compute the distance between the estimated distribution (^$\hat{\mathcal{F}}_{B}$) and the true distribution ($\mathcal{F}_{B}$) by summing the absolute difference of each bin’s value. Formally,
^$\mid\hat{\mathcal{F}}_{B}-\mathcal{F}_{B}\mid\coloneq\sum_{i=1}^{50}\mid\hat{\mathcal{F}}_{B_{i}}-\mathcal{F}_{B_{i}}\mid,$
where ^$\hat{\mathcal{F}}_{B_{i}}$ and $\mathcal{F}_{B_{i}}$ are the $i$-th bin’s value of the estimated and true distributions, respectively.  Although this approach cannot capture the property of the circular data, it remains the most viable metric for comparing circular distributions (Fisher, 2000; Mardia and Jupp, 2009). Under a more relevant approach, the performance of OGPM for the circular domain could be even better.
For mean estimation, we compute the absolute difference between the estimated and true mean, i.e. $|\hat{\mu}-\mu|$, where $\hat{\mu}$ is the estimated mean and $\mu$ is the true mean in the classical domain or the circular domain. In the classical domain, the true mean is $\mu=\frac{1}{n}\sum_{i=1}^{n}x_{i}$. In the circular domain, the mean is computed by the circular mean formula (Fisher, 2000; Mardia and Jupp, 2009):
$\mu=\text{atan2}\left(\frac{1}{n}\sum_{i=1}^{n}\sin x_{i},\frac{1}{n}\sum_{i=1}^{n}\cos x_{i}\right).$
We repeat the experiments $500$ times for stable results and report the average error.
7.2.2. Distribution Estimation
Figure 16 shows the comparison of the errors of distribution estimation in the classical and circular domains. We can see that OGPM outperforms SW and PM with smaller errors in both types of domains.  In the classical domain, OGPM’s error decreases faster when $\varepsilon$ increases above $3$, and SW-C decreases slower than PM-C, consistent with the expected error comparison.  In the circular domain, OGPM’s error is significantly lower than SW-C and PM-C, despite the limitation of the distance metric used for circular distributions. We also observe that SW-C performs better than PM-C in this domain. This is because SW has higher sampling probabilities for both the central piece and other pieces, making it sample the true value more frequently when $\varepsilon$ is large in practice in a large-size domain, despite the large expected error theoretically.  Statistically, OGPM’s distribution error is $93.5\%$ of PM-C and $86.7\%$ of SW-C in the classical domain, and $72.2\%$ of PM-C and $84.0\%$ of SW-C in the circular domain.
7.2.3. Mean Estimation
Figure 17 shows the comparison of the errors of mean estimation in the classical and circular domains. OGPM consistently outperforms other mechanisms in both types of domains, with significantly lower errors.  In the classical domain, we also compare with OGPM-U, which is designed for unbiased mean estimation, but it has an enlarged output domain, leading to a higher distance between the sensitive and randomized data. We can see that such larger output domain results in higher mean estimation errors when $\varepsilon$ is small, but OGPM-U outperforms SW when $\varepsilon\geq 5.5$.  In the circular domain, OGPM’s mean estimation error is negligible. We also observe that SW-C performs better than PM-C in this large-size domain.  Statistically, OGPM’s mean estimation error is $66.2\%$ of PM-C and $55.4\%$ of SW-C in the classical domain. In the circular domain, OGPM’s mean estimation error is merely $2.3\%$ of PM-C and $3.6\%$ of SW-C.
(a) Classical domain.
(b) Circular domain.
Figure 16. Comparison of distribution estimation error.
(a) Classical domain.
(b) Circular domain.
Figure 17. Comparison of mean estimation error.
8. Related Work
This paper focuses on the optimal mechanism for collecting numerical data with bounded domains under LDP, related to numerical data collection and the optimality of DP mechanisms.
Numerical Data Collection under LDP
Classical noise-adding mechanisms such as the Laplace and Gauss mechanisms (Dwork et al., 2006; Dwork and Roth, 2014) add randomly sampled noise from a distribution to the data to achieve LDP. However, they generate outputs in an unbounded domain due to the unbounded noise distributions, rendering them unsuitable for applications requiring bounded domains (Wang et al., 2019).
For bounded data domains, the basic idea is to sample the output from a carefully designed distribution on the bounded domain. Duchi et al. randomize any data in $[-1,1)$ to two discrete values $y\in\{-C_{\varepsilon},C_{\varepsilon}\}$ (Duchi et al., 2016), where $C_{\varepsilon}$ is a constant depending on the privacy level $\varepsilon$. This binary-output mechanism exhibits a large randomization error as the output space is too coarse.  PM (Wang et al., 2019) extends the binary output to a continuous output in $[-C_{\varepsilon},C_{\varepsilon})$. It designs a piecewise-based mechanism to sample the output, which uses different sampling intervals for different $x$, achieving a lower randomization error. Both PM and later SW (Li et al., 2020) have shown the data utility advantage of TPM in numerical data collection with bounded domains under LDP. PTT (Ma et al., 2024) discusses the optimality of TPM. It shows that there exist TPM instantiations that yield optimal data utility, but it does not provide the closed-form mechanism.  TPM is a special case of GPM using $3$-piece distributions and a specific form for each piece. Meanwhile, existing TPM instantiations are designed for specific error metrics and only classical domains.
Applications of TPM. A natural application of TPM is in high-dimensional numerical data. This includes scenarios where the sensitive data may be a high-dimensional vector (Wang et al., 2019; Zhang et al., 2023), an infinite data stream (Ren et al., 2022), or matrixes (Wang et al., 2023).  Another application area is federated learning. TPM ensures LDP in the $[0,1)$ domain, which is commonly used as the normalized domain in model training. TPM avoids the data clipping required by the noise-adding mechanisms (Abadi et al., 2016; Liu et al., 2023; Fu et al., 2024).  Another typical application is in sensors, where the data is bounded by the sensor’s physical nature. This paper can replace the existing TPM to achieve better utility.
A recent work aims to design other types of bounded distributions besides TPM to achieve (L)DP (Zhang et al., 2024). They tailor sin functions and quartic functions on the bounded domain to satisfy the DP constraint. From their experimental results, the piecewise-based design is the best choice among their six instantiations for bounded domains under DP. This also indicates the advantages of piecewise-based mechanisms in the bounded domain.
Optimality of DP Mechanisms
Achieving optimal data utility is a common concern across all DP mechanisms. The staircase mechanism showed that the Laplace mechanism does not generate optimal noise (Soria-Comas and Domingo-Ferrer, 2013; Geng and Viswanath, 2014). It samples noise from a staircase distribution, which has been shown to achieve lower error compared to the Laplace mechanism.  Besides optimality in concrete error values (which is the focus of this paper), another widely focused concept is asymptotic optimality (Cai et al., 2019; Ma et al., 2024), which studies optimal asymptotic bounds of a mechanism or a statistical estimation. Appendix B.2 discusses more detailed optimalities.
Beyond the optimality of a single DP procedure, the optimality of multifold compositions such as iterative training under DP, is studied by advanced compositions (Murtagh and Vadhan, 2018; Kairouz et al., 2017; Dong et al., 2019, 2020). In high-dimensional settings, the sensitivity set across dimensions also influences the optimality (Xiao et al., 2023), because it affects the choice of the privacy parameter. These works either focus on different privacy constraints or discuss optimality beyond a single DP procedure, making them orthogonal to ours.
9. Conclusions
This paper presents the optimal piecewise-based mechanism for collecting numerical data with bounded domains under LDP. To find the optimal mechanism among all possible piecewise mechanisms, we generalize the existing $3$-piece mechanism to an $m$-piece mechanism with the most general form. We proposed a framework that combines analytical proofs and off-the-shelf optimization solvers to find the optimal mechanism. Our results include the closed-form optimal piecewise mechanisms for both the classical and circular domain. Theoretical and experimental evaluations confirm the advantages of our mechanisms over existing mechanisms.
Appendix A Proofs
Notations. Table 3 provides the notations used throughout this paper.
A.1. Proof of Lemma 3.2
Proof.
We prove the inner $\max_{x}$ problem has a closed form. According to the definition of $\mathcal{P}_{\mathcal{M}(x)}$, it is
$\max_{x}\int_{\tilde{\mathcal{D}}}\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}\mathrm{d}y=\max_{x}\sum_{i=1}^{m}p_{i}\int_{l_{i}}^{r_{i}}\mathcal{L}(y,x)\mathrm{d}y.$
Denote $f_{i}(x)\coloneq\int_{l_{i}}^{r_{i}}\mathcal{L}(y,x)\mathrm{d}y$, where $\mathcal{L}(y,x)=|y-x|^{p}$. First, we prove that $f_{i}(x)$ is a convex function w.r.t. $x$.  Based on the relationship between $x$ and $[l_{i},r_{i})$, the value of $x$ is split into three cases: (i) $x\in[a,l_{i})$, (ii) $x\in[l_{i},r_{i})$, and (iii) $x\in[r_{i},b)$. We prove the second derivative of $f_{i}(x)$ w.r.t. $x$ is non-negative in each case, thus $f_{i}(x)$ is convex.
Case (i): $x\in[a,l_{i})$. The integral is:
$\begin{split}f_{i}(x)&=\int_{l_{i}}^{r_{i}}|y-x|^{p}\mathrm{d}y=\int_{l_{i}}^{r_{i}}(y-x)^{p}\mathrm{d}y\\ &=\frac{(r_{i}-x)^{p+1}-(l_{i}-x)^{p+1}}{p+1}.\end{split}$
The second derivative w.r.t. $x$ is
$\frac{\partial^{2}}{\partial x^{2}}f_{i}(x)=p(r_{i}-x)^{p-1}-p(l_{i}-x)^{p-1}\geq 0.$
The inequality holds because $(r_{i}-x)^{p-1}\geq(l_{i}-x)^{p-1}$ for $x\in[a,l_{i})$.
Case (ii): $x\in[l_{i},r_{i})$. The integral is
$\begin{split}f_{i}(x)&=\int_{l_{i}}^{r_{i}}|y-x|^{p}\mathrm{d}y\\ &=\int_{l_{i}}^{x}(x-y)^{p}\mathrm{d}y+\int_{x}^{r_{i}}(y-x)^{p}\mathrm{d}y\\ &=\frac{(x-l_{i})^{p+1}}{p+1}+\frac{(r_{i}-x)^{p+1}}{p+1}.\end{split}$
The second derivative w.r.t. $x$ is
$\begin{split}\frac{\partial^{2}}{\partial x^{2}}f_{i}(x)&=p(x-l_{i})^{p-1}+p(r_{i}-x)^{p-1}\geq 0.\end{split}$
The inequality holds because both $(x-l_{i})^{p-1}$ and $(r_{i}-x)^{p-1}$ are non-negative for $x\in[l_{i},r_{i})$.
Case (iii): $x\in[r_{i},b)$. The integral is
$\begin{split}f_{i}(x)&=\int_{l_{i}}^{r_{i}}|y-x|^{p}\mathrm{d}y=\int_{l_{i}}^{r_{i}}(x-y)^{p}\mathrm{d}y\\ &=\frac{(x-l_{i})^{p+1}-(x-r_{i})^{p+1}}{p+1}.\end{split}$
The second derivative w.r.t. $x$ is
$\begin{split}\frac{\partial^{2}}{\partial x^{2}}f_{i}(x)&=p(x-l_{i})^{p-1}-p(x-r_{i})^{p-1}\geq 0.\end{split}$
The inequality holds because $(x-l_{i})^{p-1}\geq(x-r_{i})^{p-1}$ for $x\in[r_{i},b)$.
The above three cases show that the second derivative of $f_{i}(x)$ w.r.t. $x\in[a,b)$ is non-negative. Thus, the non-negative weighted sum $\sum_{i=1}^{m}p_{i}f_{i}(x)$ is also convex w.r.t. $x$ (Boyd and Vandenberghe, 2004).  According to the Bauer maximum principle (Wikipedia, 2024): any function that is convex attains its maximum at some extreme points of set. This means that the optimal $x$ is achieved at the endpoints of $x\in\mathcal{D}$, i.e. $x=a$ or $x=b$. Therefore, we have
$\max_{x}\int_{\tilde{\mathcal{D}}}\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}\mathrm{d}y=\max_{\{a,b\}}\int_{\tilde{\mathcal{D}}}\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}\mathrm{d}y,$
which completes the proof. ∎
Remark: This lemma can be empirically validated by the whole-domain error plots in Figure 9 and Figure 11, where the maximum of the whole-domain error is achieved at the endpoints.
Table 3. Notations
Symbol Description
$x$ Sensitive input (from raw data)
$y$ Randomized output
$\mathcal{D}$ Input domain
$\tilde{\mathcal{D}}$ Output domain
$pdf[\cdot]$ Probability density function
$\mathcal{P}_{\mathcal{M}(x)}$ Probability density function of $\mathcal{M}(x)$
$p_{\varepsilon}$ Sampling probability w.r.t. $\varepsilon$
$[l_{x,\varepsilon},r_{x,\varepsilon})$ Sampling interval w.r.t. $x$ and $\varepsilon$
A.2. Proof of Lemma 3.3
Proof.
The optimal $(m+1)$-piecewise mechanism and the optimal $m$-piecewise mechanism may superficially differ due to the extra piece. Thus, we define a piece-merging operation to merge the redundant pieces. We will show that if the optimal $(m+1)$-piecewise mechanism is the same as the optimal $m$-piecewise mechanism after merging redundant pieces, then increasing $m$ does not decrease the optimal error, i.e. the optimal piece number is $m$.
Assume the optimal $m$-piecewise mechanism is determined by the tuple set
$S_{m}=\{(p_{i},l_{i},r_{i}):i\in[m]\}.$
To merge redundant pieces, we define a piece-merging operation:
$(p_{i},l_{i},r_{i})\uplus(p_{j},l_{j},r_{j})\coloneq\begin{dcases}(p_{i},l_{i},r_{j})\quad\text{if}\ p_{i}=p_{j}\ \text{and}\ i+1=j,\\ \{(p_{i},l_{i},r_{i}),(p_{j},l_{j},r_{j})\}\quad\text{otherwise}.\end{dcases}$
Because the optimal ($m+1$)-piecewise mechanism is the same as the $m$-piecewise mechanism, it follows that
$\uplus_{i,j=1}^{m+1}S_{m+1}=\uplus_{i,j=1}^{m}S_{m},$
where $\uplus_{i,j=1}^{m}S_{m}$ merges all consecutive pieces with the same $p$. Denote the merged optimal $m$-piecewise mechanism as $\uplus_{i,j=1}^{m}S_{m}\coloneq S_{m}^{\uplus}$ and the piece number as $|S_{m}^{\uplus}|=m^{*}$. Because both sides of the above equation are optimal, this means that if $(p_{k},l_{k},r_{k})$ is an arbitrary piece in the optimal ($m+1$)-piecewise mechanism, then merging it with $S_{m}^{\uplus}$ remains $S_{m}^{\uplus}$, i.e.
$(p_{k},l_{k},r_{k})\uplus_{i=1}^{m^{*}}S_{m}^{\uplus}=S_{m}^{\uplus}.$
This premise indicates that there does not exist a piece $(p_{k},l_{k},r_{k})$ besides $S_{m}^{\uplus}$ lowers the error.
Without loss of generality, we can consider the optimal $(m+2)$-piecewise mechanism, which allows an extra piece besides the optimal $(m+1)$-piecewise mechanism. We claim that the extra piece is still captured by $S_{m}^{\uplus}$.  The key insight is: adding an extra optimal piece to the optimal $(m+1)$-piecewise mechanism is the same as adding it to the optimal $m$-piecewise mechanism, because the optimal $(m+1)$-piecewise mechanism is the same as the optimal $m$-piecewise mechanism.
Since adding an extra optimal piece to the optimal $m$-piecewise mechanism remains $S_{m}^{\uplus}$, then for any $k\in[m+2]$, merging piece $(p_{k},l_{k},r_{k})$ in the optimal ($m+2$)-piecewise mechanism remains an $m$-piecewise mechanism $S_{m}^{\uplus}$.
For $m+3$ or more, it follows the same logic. Adding an arbitrary piece is equivalent to adding it to the optimal $(m+2)$-piecewise mechanism, which is the same as adding it to the optimal $m$-piecewise mechanism. Thus, the optimal $m$-piecewise mechanism is the same as the optimal $(m+3)$-piecewise mechanism after merging redundant pieces, and so on. ∎
Remark: Intuitively, the extra pieces (of the optimal $(m+1)$-piecewise mechanism and beyond) is similar to the redundant variables in optimization theory: adding more non-negative variables to a minimization objective does not decrease the optimal value. Here the error from each piece is a variable, and it is non-negative, which leads to the same conclusion: adding more pieces (one or more) to the optimal $m$-piecewise mechanism does not decrease the optimal error.
This lemma means that we can determine the optimal $m$-piecewise mechanism for $m=1,2,\ldots$, until the optimal $(m+1)$-piecewise mechanism is identical to the optimal $m$-piecewise mechanism for all $x$ and $\varepsilon$. This statement can be empirically validated by attempting to find counterexamples using larger $m$ than the optimal $m$. The source code of our framework provides this validation.
A.3. Proof of Theorem 3.6
Proof.
Privacy invariant: For any input $v,v^{\prime}\in\mathcal{D}^{\prime}$ and any output $y\in\tilde{\mathcal{D}}^{\prime}$:
$\frac{pdf\left[\mathcal{M}^{\prime}(v)=y\right]}{pdf\left[\mathcal{M}^{\prime}(v^{\prime})=y\right]}\leq\frac{p}{c}\div\frac{p}{\exp(\varepsilon)c}=\exp(\varepsilon).$
Utility invariant: For any $x^{\prime}=cx+d\in\mathcal{D}^{\prime}$, we can calculate the error difference between $\mathcal{M}^{\prime}(x^{\prime})$ and $\mathcal{M}^{\prime}_{\text{bad}}(x^{\prime})$ as follows:
$\begin{split}&Err(x^{\prime},\mathcal{M}^{\prime})-Err(x^{\prime},\mathcal{M}^{\prime}_{\text{bad}})\\ =&Err(cx+d,\mathcal{M}^{\prime})-Err(cx+d,\mathcal{M}^{\prime}_{\text{bad}})\\ =&\int_{\tilde{\mathcal{D}}^{\prime}}\mathcal{L}(y,cx+d)\left(\mathcal{P}_{\mathcal{M}^{\prime}(cx+d)}-\mathcal{P}_{\mathcal{M}_{\text{bad}}^{\prime}(cx+d)}\right)\mathrm{d}y.\\ \end{split}$
Let $y_{t}=(y-d)/c$, then $\mathrm{d}y=c\mathrm{d}y_{t}$ and $y=cy_{t}+d$, where $y_{t}\in\tilde{\mathcal{D}}$. The above equation is equivalent to
$\begin{split}&Err(x^{\prime},\mathcal{M}^{\prime})-Err(x^{\prime},\mathcal{M}^{\prime}_{\text{bad}})\\ =&\int_{\tilde{\mathcal{D}}}\mathcal{L}(cy_{t}+d,cx+d)\left(\mathcal{P}_{\mathcal{M}^{\prime}(cx+d)}-\mathcal{P}_{\mathcal{M}_{\text{bad}}^{\prime}(cx+d)}\right)c\mathrm{d}y_{t}\\ =&\int_{\tilde{\mathcal{D}}}\mathcal{L}(cy_{t}+d,cx+d)\frac{1}{c}\left(\mathcal{P}_{\mathcal{M}(x)}-\mathcal{P}_{\mathcal{M}_{\text{bad}}(x)}\right)c\mathrm{d}y_{t}.\end{split}$
The last equality holds due to the definition of $\mathcal{T}:\mathcal{M}\to\mathcal{M}^{\prime}$. For $L_{p}$-similar error metric $\mathcal{L}$ (i.e. $\mathcal{L}(y,x)\coloneq|y-x|^{p}$), it follows that
$\mathcal{L}(cy_{t}+d,cx+d)=\mathcal{L}(cy_{t},cx)=c^{p}\mathcal{L}(y_{t},x).$
Thus, the above difference of $Err$ is equivalent to
$\begin{split}&Err(x^{\prime},\mathcal{M}^{\prime})-Err(x^{\prime},\mathcal{M}^{\prime}_{\text{bad}})\\ =&c^{p}\int_{\tilde{\mathcal{D}}}\mathcal{L}(y_{t},x)\left(\mathcal{M}(x)-\mathcal{M}_{\text{bad}}(x)\right)\mathrm{d}y_{t}\\ =&c^{p}\left(Err(x,\mathcal{M})-Err(x,\mathcal{M}_{\text{bad}})\right)\leq 0,\end{split}$
due to the known fact $c>0$ and $Err(x,\mathcal{M})-Err(x,\mathcal{M}_{\text{bad}})\leq 0$. ∎
Remark: Intuitively, this theorem is to prove: if $\mathcal{M}$ is a better mechanism than $\mathcal{M}_{\text{bad}}$ on $\mathcal{D}$, then it is still a better mechanism than $\mathcal{M}_{\text{bad}}$ after linearly mapping their outputs to $\mathcal{D}^{\prime}$.
A.4. Proof of Theorem 3.7
Proof.
Appendix B.4 provides the formalized procedure. Following this procedure, we show the optimal GPM under $\mathcal{D}\to\tilde{\mathcal{D}}=[0,1)\to[0,1)$ and $\mathcal{L}(y,x)=|y-x|$.
The variables in TPM are $p$, $l$, and $r$. Since it is a family of probability distributions, the normalization constraint is
$(r-l)\cdot p+(1-(r-l))\cdot p/\exp(\varepsilon)=1,$
which means the length of the central piece is
$s\coloneq r-l=\frac{\exp(\varepsilon)-p}{p(\exp(\varepsilon)-1)}.$
Without loss of generality, assume $x=0$ is the optimal point ($x=1$ is symmetric). The optimization problem for solving the optimal $p$ is
$\begin{split}&\operatorname*{arg\,min}_{p}\left(\int_{0}^{s}y\cdot p\thinspace\mathrm{d}y+\int_{s}^{1}y\cdot\frac{p}{\exp(\varepsilon)}\thinspace\mathrm{d}y\right)\\ =&\operatorname*{arg\,min}_{p}\left(\frac{s^{2}}{2}\left(p-\frac{p}{\exp(\varepsilon)}\right)+\frac{1}{2}\frac{p}{\exp(\varepsilon)}\right)\\ =&\operatorname*{arg\,min}_{p}\frac{1}{2}\left(\frac{(\exp(\varepsilon)-p)^{2}}{p(\exp(\varepsilon)-1)\exp(\varepsilon)}+\frac{p}{\exp(\varepsilon)}\right).\end{split}$
To solve the optimal $p$, we take the first-order derivative w.r.t. $p$ and set it to 00, i.e.
$\frac{\partial}{\partial p}\left(\frac{(\exp(\varepsilon)-p)^{2}}{p(\exp(\varepsilon)-1)\exp(\varepsilon)}+\frac{p}{\exp(\varepsilon)}\right)=0,$
This leads to $p=\exp(\varepsilon/2)$. Then
$s=\frac{\exp(\varepsilon/2)-1}{\exp(\varepsilon)-1}.$
Having solved $p$ and $s$, the optimal $r$ is $r=l+s$. Then the optimal $l$ is determined by
$\begin{split}\operatorname*{arg\,min}_{l}&\int_{0}^{l}(x-y)\cdot\frac{p}{\exp(\varepsilon)}\thinspace\mathrm{d}y+\int_{l}^{x}(x-y)\cdot p\thinspace\mathrm{d}y+\\ &\int_{x}^{l+s}(y-x)\cdot p\thinspace\mathrm{d}y+\int_{l+s}^{1}(y-x)\cdot\frac{p}{\exp(\varepsilon)}\thinspace\mathrm{d}y\end{split}$
This is a univariate optimization problem w.r.t. $l$. Moreover, it is a two-order polynomial w.r.t. $l$ and can be solved by analyzing the first-order and second-order derivatives. The solution is
$l=\frac{2x(p-pe^{-\varepsilon})-s(p-pe^{-\varepsilon})}{2(p-pe^{-\varepsilon})}=x-\frac{s}{2}.$
Note that when $x-s/2<0$, the above $l$ is outside the domain $[0,1)$. In this case, the optimal $l$ is $l=0$.
Relating the above deduction to Theorem 3.7, the term $s/2$ corresponds to $C$. Then the optimal $p$ is $\exp(\varepsilon/2)$, $l=x-C$, and $r=x+C$ when $x\in[C,1-C)$, which completes the proof for $\mathcal{L}(y,x)=|y-x|$. The proof for $\mathcal{L}(y,x)=|y-x|^{2}$ is similar. ∎
Remark: This proof is the same as finding the optimal $3$-piecewise distribution in domain $[0,1)$. The source code of our framework provides this validation.
(a) $x=3\pi/4$.
(b) $x=\pi$.
Figure 18. Examples of $\mathcal{L}_{\text{mod}}(y,x)$ w.r.t. $x$. Given a specific $y$, the shorter between the blue and green arcs is $\mathcal{L}_{\text{mod}}(y,x)$. $\max_{x}\mathcal{L}_{\text{mod}}(y,x)$ is achieved at $x=\pi$.
A.5. Proof of Lemma 4.1
Proof.
Note that $\mathcal{L}_{\text{mod}}(y,x)=\min\left(\mathcal{L}(y,x),\mathcal{L}(y,2\pi-x)\right)$. The key observation is that for any $y$ and $L_{p}$-similar distance metric $\mathcal{L}$, we have
$\max_{x}\mathcal{L}_{\text{mod}}(y,x)=\mathcal{L}_{\text{mod}}(y,\pi)=\mathcal{L}(y,\pi).$
Figure 18 illustrates the intuition. For any fixed $y$, it compares the length of $|y-x|$ and $|y-(2\pi-x)|$ w.r.t $x\in[0,2\pi)$. $\mathcal{L}_{\text{mod}}(y,x)$ is determined by the minimum of the two, and $\max_{x}\mathcal{L}_{\text{mod}}(y,x)$ will achieved at $x=\pi$.
The remained proof is more intuitive than the proof of Lemma 3.2, as the inner integrand $\mathcal{L}_{\text{mod}}(y,x)$ has a unique maximum at $x=\pi$ for any $y\in[l_{i},r_{i})$, making the swap of $\max_{x}$ and integration valid. Specifically, we have
$\begin{split}\max_{x}\sum_{i}^{m}p_{i}\int_{l_{i}}^{r_{i}}\mathcal{L}_{\text{mod}}(y,x)\mathrm{d}y&=\sum_{i}^{m}p_{i}\int_{l_{i}}^{r_{i}}\max_{x}\mathcal{L}_{\text{mod}}(y,x)\mathrm{d}y\\ &=\sum_{i}^{m}p_{i}\int_{l_{i}}^{r_{i}}\mathcal{L}(y,\pi)\mathrm{d}y\end{split}$
holds trivially, because all other $x$ values always result in smaller $\mathcal{L}_{\text{mod}}(y,x)$. Therefore, for any values of $p_{i}\geq 0$, $l_{i}\leq r_{i}$, the integration of $\mathcal{L}_{\text{mod}}(y,x)$ is bounded by the integration of $\mathcal{L}(y,\pi)$. ∎
A.6. Proof of Theorem 5.1
Proof.
We prove the unbiasedness of $\mathcal{M}$. The expectation of the given $\mathcal{M}$ is
$\begin{split}&\mathrm{E}[\mathcal{M}(x)]=\int_{-C}^{C+1}y\cdot\mathcal{P}_{\mathcal{M}(x)}\mathrm{d}y\\ =&\int_{-C}^{l}y\frac{p}{\exp(\varepsilon)}\mathrm{d}y+\int_{l}^{r}yp\mathrm{d}y+\int_{r}^{C+1}y\frac{p}{\exp(\varepsilon)}\mathrm{d}y\\ =&\frac{r^{2}-l^{2}}{2}\left(p-\frac{p}{\exp(\varepsilon)}\right)+\left(C+\frac{1}{2}\right)\frac{p}{\exp(\varepsilon)}.\end{split}$
Denote $s\coloneq(2C+1)(C-1)/(2C)$, which rewrites $l$ and $r$ as
$l=\frac{C+1}{2}x-\frac{C-1}{4}-\frac{s}{2},\ r=\frac{C+1}{2}x-\frac{C-1}{4}+\frac{s}{2}.$
Then the above $\mathrm{E}[\mathcal{M}(x)]$ is equivalent to
$\begin{split}&\frac{(C+1)sx-(C-1)s/2}{2}\cdot\frac{4C}{(C^{2}-1)(2C+1)}+\frac{\exp(-\varepsilon/2)}{2}\\ =&x-\frac{C-1}{2(C+1)}+\frac{\exp(-\varepsilon/2)}{2}=x,\end{split}$
leading to $\mathrm{E}[\mathcal{M}]=x$, i.e. $\mathcal{M}$ is unbiased. ∎
Appendix B Complementary Materials
B.1. Detailed Comparison with PTT (Section 2.3)
Piecewise transformation technique (PTT) (Ma et al., 2024) is a framework for $3$-piecewise mechanisms. It shows that (i) many PTT mechanisms are asymptotically optimal when used to obtain an unbiased estimator for mean of numerical data, and (ii) there is a PTT that reaches the theoretical lower bound on variance.
Under the viewpoint of this paper, type-I PTT focuses on TPM that constrains the probabilities of the central interval ($p$) and the two side intervals ($q$) as
$p=\frac{1}{2ak}\frac{\exp(\varepsilon)}{\exp(\varepsilon)-1},\quad q=\frac{\exp(\varepsilon)}{k(\exp(\varepsilon)-1)},$
where $a$ and $k$ are parameters to be determined. Type-II PTT focuses on TPM that constrains $p$ and $q$ as
$p=\frac{1}{ak}\frac{\exp(\varepsilon)}{\exp(\varepsilon)-1},\quad q=\frac{\exp(\varepsilon)}{k(\exp(\varepsilon)-1)}.$
Therefore, PTT is still a specific case of the TPM framework. Additionally, when discussing optimality of PTT, it gives a value of $a$ for type-I PTT but do not provide the optimal $k$.
B.2. Related Optimality (Section 3.1)
In this paper, the optimality of GPM is defined with respect to: (i) the worst-case $L_{p}$-similar error metric, (ii) bounded numerical domains $\mathcal{D}\to\tilde{\mathcal{D}}$ and mechanisms based on piecewise distributions, (iii) minimization of error value (not asymptotic or order-of-magnitude optimality), and (iv) without post-processing. $L_{p}$-similar error metrics are natural choices for evaluating data utility (Hardt and Talwar, 2010; Geng and Viswanath, 2014; Wang et al., 2019). Bounded numerical domains are common in real-world applications. Focusing on error values allows for more precise comparisons between different mechanisms. By excluding post-processing, we can analyze the optimality of the mechanism itself, providing a more fundamental understanding than considering the mechanism combined with specific post-processing.
Other types of optimality have been explored in the literature, particularly for variants of Laplace mechanisms. The staircase mechanism (Geng and Viswanath, 2014) adopts the same utility model without prior knowledge or post-processing as this paper. It claims optimality under specific assumptions, one of which is that a staircase (piecewise) distribution can achieve the optimal error. The mechanism demonstrates better $L_{1}$-error performance than the Laplace mechanism on $\tilde{\mathcal{D}}=(-\infty,\infty)$, and its asymptotic optimality has been formally proven.  Universal optimality is another type of optimality, defined from the perspective of a user’s prior knowledge and post-processing ability (Ghosh et al., 2009). In this utility model, the user observes the output of the mechanism and selects another value based on the output and their prior knowledge, i.e. under a Bayesian utility framework. Formally, if the user’s prior is denoted as $p_{i}$ on the data domain $i\in N$ (i.e. a discrete domain) and the user’s post-processing is represented as a remap $z_{i,j}$ that reinterprets the output of the mechanism (on the sensitive value $i$) to $j$, then the utility model is
$Err(i)=\sum_{i\in N}p_{i}\sum_{j\in N}z_{i,j}\cdot\mathcal{L}(i,j).$
This utility model incorporates the user’s prior knowledge and post-processing ability. A mechanism is called universally optimal if, for any prior $p_{i}$, there exists an optimal remap $z_{i,j}$. Under this utility model, it was proven that the truncated geometric mechanism (a discretized version of the Laplace mechanism) can achieve universal optimality for count queries111111 This is in the centralized DP setting, where the data curator holds the dataset and uses one mechanism. and a legal error metric $\mathcal{L}(i,j)$. Such universal optimality was shown to be unachievable for more complex queries (Brenner and Nissim, 2014). Under the same utility model, the universal optimality was extended to the truncated Laplace mechanism for a bounded numerical domain $\mathcal{D}=[0,1]$ by approximating the geometric mechanism with the Laplace mechanism and post-processing (Fernandes et al., 2021).
These results do not hold in our utility model, i.e. utility model without prior and post-processing. Figure 13 has shown that OGPM generally has a smaller error than the truncated Laplace mechanism, indicating the sub-optimality of the truncated Laplace mechanism in the absence of prior and post-processing.
B.3. Directions for Analytically Proving Optimal $m=3$ (Section 3.2.3)
This appendix outlines two potential directions for analytically proving that the optimal $m$ is $3$, along with the challenges associated with each approach.
Mathematically, finding the optimal $m$-piecewise mechanism is equivalent to identifying the optimal $m$-piecewise distribution under an $L_{p}$-similar error metric. It is seemingly true that the optimal $m$ is $3$: if the optimal $m$-piecewise distribution is not $3$ but $4$ or more, we can always shift the probability mass from the two side intervals (i.e. other pieces) to the central interval, thereby reducing the error. At the very least, the following fact holds:
Fact B.1.
The optimal $m$-piecewise distribution has a strict staircase shape, i.e. the probability density of the central interval is greater than that of the two side intervals.
(a) Non-staircase distribution.
(b) Staircase distribution.
Figure 19. A non-staircase distribution (left) can always be shifted into a staircase distribution (right) by moving some pieces closer to $x$, which reduces the error.
Figure 19 illustrates this fact. Moving pieces while keeping their probabilities unchanged clearly maintains both the $\varepsilon$-LDP constraint and the probability normalization constraint. This observation reduces the problem to proving that a $3$-staircase distribution can achieve the same optimal error as a $4$-staircase distribution under the $\varepsilon$-LDP and probability normalization constraints.
Direction 1: If we can further move the green piece in Figure 19(b) “into” the red central piece while keeping the probabilities of the red and blue pieces unchanged, i.e. transform it into a $3$-staircase distribution while ensuring a decrease in the error, then we can prove that the optimal $m$ is $3$. However, this is challenging, as it breaks the probability normalization constraint, requiring adjustments to the probabilities of each piece to satisfy the $\varepsilon$-LDP constraint. The difficulty lies in ensuring that these adjustments will indeed decrease the error.
Direction 2: Another approach is to formulate the problems for $3$-staircase and $4$-staircase distributions as two constrained optimization problems. The goal would be to prove that the optimal error of the $3$-staircase distribution is equivalent to that of the $4$-staircase distribution. Ideally, these two multi-variable optimization problems could be solved analytically, resulting in two closed-form error expressions w.r.t. $x$ and $\varepsilon$, thereby completing the proof for any $x$ and $\varepsilon$. This direction aligns with our framework. However, the challenge lies in the complexity of solving such multi-variable optimization problems analytically. This is why we rely on an off-the-shelf optimization solver, which, while effective, only provides numerical solutions for specific $x$ and $\varepsilon$ values.
B.4. Analytical Deduction for the Optimal GPM (Section 3.2.3)
If the optimal GPM under $\mathcal{L}$ is proved to be TPM, then the closed-form optimal can be derived by deduction.
Denote $\mathcal{D}=[a,b)$ and $\tilde{\mathcal{D}}=[\tilde{a},\tilde{b})$. Notice that the normalization constraint of probability is
$(r-l)\cdot p+[(b-a)-(r-l)]\cdot p/\exp(\varepsilon)=1.$
This means the central interval length is
$s\coloneq r-l=\frac{1}{p(1-\exp(-\varepsilon))}-\frac{\tilde{b}-\tilde{a}}{\exp(\varepsilon)-1}.$
If the minimal worst-case error is achieved at $x=a$ in Lemma 3.2, then solving for the optimal $p$ is reduced to
$\operatorname*{arg\,min}_{p}\left(\int_{\tilde{a}}^{s}\mathcal{L}(y,a)p\thinspace\mathrm{d}y+\int_{s}^{\tilde{b}}\mathcal{L}(y,a)\frac{p}{\exp(\varepsilon)}\mathrm{d}y\right),$
which is a univariate optimization problem w.r.t. $p$ and can be solved analytically. With the solved $p$, Formulation (3) is also reduced to a univariate optimization problem w.r.t. $l$:
$\operatorname*{arg\,min}_{l}\left(\int_{\tilde{a}}^{l}P_{1}\thinspace\mathrm{d}y+\int_{l}^{l+s}P_{2}\thinspace\mathrm{d}y+\int_{l+s}^{\tilde{b}}P_{1}\thinspace\mathrm{d}y\right),$
where $P_{1}=\mathcal{L}(y,x)p$ and $P_{2}=\mathcal{L}(y,x)p/\exp(\varepsilon)$. This univariate optimization problem solves the optimal $l$, and the optimal $r$ is $r=l+s$. Note that $l$ and $r$ should be restricted in $[\tilde{a},\tilde{b})$ when analyzing the first-order derivative.
B.5. MSE of the Optimal GPM (Section 3.3)
Denote $p_{\varepsilon}$ and $C$ as the same as the instantiations of $\mathcal{M}$ in Theorem 3.7. The MSE of $\mathcal{M}$ is
  1. (1)
    If $x\in[0,C)$:
    $\frac{p_{\varepsilon}}{3}\left((2C-x)^{3}+x^{3}\right)+\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left((1-x)^{3}-(2C-x)^{3}\right).$
  2. (2)
    If $x\in[C,1-C)$:
    $\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left(-2C^{3}+3x^{2}-3x+1\right)+\frac{p_{\varepsilon}}{3}\left(2C^{3}\right).$
  3. (3)
    If $x\in[1-C,1)$:
    $\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left((1-2C-x)^{3}+x^{3}\right)+\frac{p_{\varepsilon}}{3}\left((1-x)^{3}-(1-2C-x)^{3}\right).$
For example, when $x=0$, the MSE of $\mathcal{M}$ is
$MSE(\mathcal{M}(0),0)=\frac{p_{\varepsilon}}{3}\left(8C^{3}\right)+\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left(1-8C^{3}\right).$
Setting $\varepsilon=1$ results in $MSE(\mathcal{M}(0),0)=0.22$. As a comparison, SW (Li et al., 2020), which also designed for $\mathcal{D}=[0,1)$, has an MSE of $0.29$ at $x=0$.
(a) Privacy parameter $\varepsilon=0.4$.
(b) Privacy parameter $\varepsilon=0.8$.
Figure 20. Whole-domain error comparison in the classical domain with error metric $\mathcal{L}=|y-x|$.
B.6. Optimal Assignment of Privacy Parameter and an Example (Section 6)
The objective function to minimize the total error in 2D polar coordinates is
$\min_{\varepsilon_{1},\varepsilon_{2}}Err_{\text{1,wor}}(\varepsilon_{1})+Err_{\text{2,wor}}(\varepsilon_{2}),$
where $Err_{\text{1,wor}}(\varepsilon_{1})$ and $Err_{\text{2,wor}}(\varepsilon_{2})$ are the worst-case errors of the classical domain and the circular domain, respectively.
Without loss of generality, we can assume the polar coordinate data is in $[0,1)\times[0,2\pi)$ and $\mathcal{L}=|y_{1}-x_{1}|^{2}$, then $Err_{\text{1,wor}}(\varepsilon_{1})$ and $Err_{\text{2,wor}}(\varepsilon_{2})$ are already given by our MSE analysis. However, the above optimization problem as a function of $\varepsilon_{1}$ and $\varepsilon_{2}$ is generally non-linear, thus hard to be analytically solved. Therefore, a simple and practical way to find the optimal $\varepsilon_{1}$ and $\varepsilon_{2}$ is numerical testing.
Under distance metric $\mathcal{L}(y,x)=|y-x|^{2}$, the worst-case error of
the classical domain $[0,1)$ is achieved at $x=0$. Therefore, $Err_{\text{1,wor}}$ equals to $MSE(\mathcal{M}(0),0)$ calculated before. For the circular domain $[0,2\pi)$, the worst-case error $Err_{\text{2,wor}}$ is stated in Theorem 4.4.  If $\varepsilon=1+2\pi$ and we assign $\varepsilon_{1}$ to the classical domain and $\varepsilon_{2}=\varepsilon-\varepsilon_{1}$ to the circular domain, then the total error is plotted in the right figure. The optimal assignment is $\varepsilon_{1}=1.32$ and $\varepsilon_{2}=5.69$.
From the curve of the total error, we can see that $\varepsilon_{2}$ affects the total error more than $\varepsilon_{1}$. Even if $\varepsilon_{1}$ is set to 00, the total error is not significantly affected, and it is still is significantly smaller than the case of $\varepsilon_{2}=0$. This is because the circular domain has a larger range than the classical domain, thus the error of the circular domain is more sensitive to the privacy parameter.
Note that the optimality for the polar coordinate data is under the specific error metric $\mathcal{L}_{\text{2D}}\coloneq\mathcal{L}(y_{1},x_{1})+\mathcal{L}_{\text{mod}}(y_{2},x_{2})$. If the error metric differs, the optimal error might not be preserved.
B.7. Comparison under Small $\varepsilon$ (Section 7.1.1)
Figure 20 presents the whole-domain error comparison of OGPM, PM-C, and SW-C under smaller $\varepsilon$ values, specifically $\varepsilon=0.4$ and $\varepsilon=0.8$. In these scenarios, all three mechanisms approach the uniform distribution more closely compared to cases with larger $\varepsilon$. Consequently, their errors are also more similar to each other. Statistically, when $\varepsilon=0.4$, the error of OGPM is at most $0.008$ smaller than that of PM-C and SW-C. For $\varepsilon=0.8$, the error of OGPM is at most $0.015$ smaller than that of PM-C and SW-C.
B.8. Expected Error of the B-Laplace Mechanism (Section 7.1.3)
The B-Laplace mechanism redefines a Laplace-shaped distribution on a bounded domain as the perturbation mechanism. For the domain $\mathcal{D}\to\tilde{\mathcal{D}}=[0,1)\to[0,1)$, the B-Laplace mechanism is defined as follows:
Definition B.1 (Bounded Laplace Mechanism, adapted from (Holohan et al., 2020)).
The B-Laplace mechanism $\mathcal{M}(x):[0,1)\to[0,1)$ is given by the probability density function (PDF) as follows:
$pdf[\mathcal{M}(x)=y]=\frac{1}{C_{y}}\cdot\frac{1}{2b}\exp\left(-\frac{|y-x|}{b}\right)\quad\forall y\in[0,1),$
where $b$ is the scale parameter, and $C_{y}=\int_{0}^{1}\frac{1}{2b}\exp\left(-\frac{|y-x|}{b}\right)\mathrm{d}x$ is the normalization constant.
According to Theorem 3.5 and Corollary 4.5 in (Holohan et al., 2020), the B-Laplace mechanism satisfies $\varepsilon$-LDP whenever $b\geq 1/\varepsilon$. Using the best scale parameter $b=1/\varepsilon$, the normalization constant becomes $C_{y}=(1-\exp(-\varepsilon))/2$. We can compute the expected $L_{1}$ error of the B-Laplace mechanism as follows (this computation is not included in (Holohan et al., 2020)):
[…]
The above integral can be numerically computed using the Python library function scipy.stats.laplace.expect() or analytically solved. The final result for the expected error is
$\frac{2-(1+\varepsilon x)e^{-\varepsilon x}-(1+\varepsilon(1-x))e^{-\varepsilon(1-x)}}{\varepsilon(1-e^{-\varepsilon})},$
which is a closed-form expression w.r.t. $x$ and $\varepsilon$.

### v2

@@
 This paper investigates the optimal design of piecewise-based mechanisms to maximize data utility under LDP.
-We demonstrate that existing piecewise-based mechanisms are heuristic forms of the $3$-piecewise mechanism, which is far from enough to study optimality.
+We demonstrate that existing piecewise-based mechanisms are heuristic instances of the $3$-piecewise mechanism, which is far from enough to study optimality.
 We generalize the $3$-piecewise mechanism to its most general form, i.e.
@@
 These types of data are crucial for third-party platforms to provide personalized services.
-However, collecting such data involves particular privacy concerns, as the third-party collectors are potentially untrusted and the data often contain sensitive information.
+However, collecting such data involves particular privacy concerns, as third-party collectors are potentially untrusted and the data often contain sensitive information.
 Simple anonymization techniques (Yang et al., 2013; Niu et al., 2014) have been proven insufficient to prevent privacy leakage (Wikipedia contributors, 2023; Narayanan and Shmatikov, 2008; Yu et al., 2007).
@@
 The optimality of piecewise-based mechanisms remains a challenging problem.
-We will see that existing instantiations are heuristic forms of the $3$-piecewise mechanism (TPM).
+We will see that the existing instantiations are heuristic forms of the $3$-piecewise mechanism (TPM).
 As a special case with $3$ pieces and pre-defined forms of pieces, TPM is far from enough to study the optimality of piecewise-based mechanisms.
@@
 The results prove and validate our mechanisms’ advantages over existing mechanisms.
-Codes are available at https://github.com/ZhengYeah/Optimal-GPM.
+The codes are available at https://github.com/ZhengYeah/Optimal-GPM.
 Structure.
@@
 A lower privacy parameter $\varepsilon\in[0,+\infty)$ means higher privacy.
-For example, $\varepsilon=0$ requires $\mathcal{M}$ maps two arbitrary inputs to any output $y$ with the same probability, thus $y$ contains no distribution information about $x$, making any hypothesis-testing method to infer the sensitive $x$ powerless even with known $\mathcal{M}$.
+For example, $\varepsilon=0$ requires that $\mathcal{M}$ maps two arbitrary inputs to any output $y$ with the same probability, thus $y$ contains no distribution information about $x$, making any hypothesis-testing method to infer the sensitive $x$ powerless even with known $\mathcal{M}$.
 2.3. Piecewise-based Mechanisms
@@
 When the input domain $\mathcal{D}$ is both continuous and bounded, the state-of-the-art mechanisms to achieve $\varepsilon$-LDP are piecewise-based mechanisms.
-We summarize these mechanisms as heuristic forms of the following definition.
+We summarize these mechanisms as heuristic instances of the following definition.
 Definition 2.2.
@@
 None of the existing instantiations provided closed forms for the optimal data utility.
-Meanwhile, they also assume invariable length $r_{x,\varepsilon}-l_{x,\varepsilon}$ of the central piece for all $x$, and symmetric probability $p_{\varepsilon}/\exp(\varepsilon)$ for the remaining two pieces.
+Meanwhile, they also assume an invariable length $r_{x,\varepsilon}-l_{x,\varepsilon}$ of the central piece for all $x$, and symmetric probability $p_{\varepsilon}/\exp(\varepsilon)$ for the remaining two pieces.
 However, a general-form piecewise-based mechanism can have more pieces, unfixed piece lengths, and asymmetric probabilities, potentially improving data utility.
@@
 We detail them as follows.
-  • (Solving difficulty) Formulation (2) is a min-max problem, and the integrand $\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}$ is non-linear; it is a non-convex problem whose global optimal hard to solve.
+  • (Solving difficulty) Formulation (2) is a min-max problem, and the integrand $\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}$ is non-linear.
+It is a non-convex problem whose global optimal hard to solve.
   • (Optimal $m$) The solved optimal is only for $m$-GPM given $m$. It is necessary to find the optimal piece number $m$.
@@
 PM (Wang et al., 2019) on $\mathcal{D}=[-1,1)$, we can transform the optimal GPM to $[-1,1)$ and compare their MSE.
-Detailed comparison with them is presented in the evaluation section.
+Detailed comparisons with them are presented in the evaluation section.
 4. Optimal GPM for Circular Domain
@@
 Figure 5 summarizes the above two reductions to solve the optimal $p_{i}$, $l_{i,x}^{\text{mod}}$ and $r_{i,x}^{\text{mod}}$ in the circular domain.
-Figure 5. Reduced forms of solving the optimal $p_{i}$, $l_{i,x}^{\text{mod}}$ and $r_{i,x}^{\text{mod}}$. Optimizations under $\mathcal{L}_{\text{mod}}$ can be reduced to those under $\mathcal{L}$.
+Figure 5.
+Reduced forms of solving the optimal $p_{i}$, $l_{i,x}^{\text{mod}}$ and $r_{i,x}^{\text{mod}}$.
+Optimizations under circular distance $\mathcal{L}_{\text{mod}}$ can be reduced to those under linear distance $\mathcal{L}$.
 4.2. Closed Form for Circular Domain
@@
 Therefore, the MSE of the optimal GPM in Theorem 4.2 has an identical value for all $x$ in the circular domain, which is
-$\begin{split}&MSE(\mathcal{M}(x),x)=MSE(\mathcal{M}(\pi),\pi)\\ =&2\left(\int_{0}^{l_{\pi,\varepsilon}^{\textup{mod}}}(y-\pi)^{2}\frac{p_{\varepsilon}}{\exp(\varepsilon)}\mathrm{d}y+\int_{l_{\pi,\varepsilon}^{\textup{mod}}}^{\pi}(y-\pi)^{2}p_{\varepsilon}\mathrm{d}y\right).\end{split}$
+$\begin{split}&\mathrm{MSE}[\mathcal{M}(x)]=\mathrm{MSE}[\mathcal{M}(\pi)]\\ =&2\left(\int_{0}^{l_{\pi,\varepsilon}^{\textup{mod}}}(y-\pi)^{2}\frac{p_{\varepsilon}}{\exp(\varepsilon)}\mathrm{d}y+\int_{l_{\pi,\varepsilon}^{\textup{mod}}}^{\pi}(y-\pi)^{2}p_{\varepsilon}\mathrm{d}y\right).\end{split}$
 Calculating the above integral, we can obtain the MSE of the optimal GPM in the circular domain.
@@
 The optimal GPM in Theorem 4.2 has an identical MSE for all $x$ in the circular domain, which is
-$MSE(\mathcal{M}(x),x)=\frac{2}{3}\left((\pi^{3}-C^{3})\frac{p_{\varepsilon}}{\exp(\varepsilon)}+C^{3}p_{\varepsilon}\right),$
+$\mathrm{MSE}[\mathcal{M}(x)]=\frac{2}{3}\left((\pi^{3}-C^{3})\frac{p_{\varepsilon}}{\exp(\varepsilon)}+C^{3}p_{\varepsilon}\right),$
 where $C=\pi(\exp(\varepsilon/2)-1)/(\exp(\varepsilon)-1)$.
 Proof.
-We have shown that the MSE at $x$ is the same as that at $\pi$. Then the calculation is straightforward by plugging in the values of $p_{\varepsilon}$ and $l_{\pi,\varepsilon}^{\textup{mod}}$. ∎
+We have shown that the MSE at $x$ is the same as that at $\pi$.
+Then the calculation is straightforward by plugging in the values of $p_{\varepsilon}$ and $l_{\pi,\varepsilon}^{\textup{mod}}$ in Theorem 4.2.
+∎
 If we do not consider $\mathcal{L}_{\text{mod}}$ and directly apply the optimal GPM in the classical domain (or SW and PM mechanisms) to the circular domain, i.e.
@@
 The accuracy of this estimator is measured by $|\hat{\mu}-\mu|$, where $\mu$ is the true mean of $\mathcal{X}$.
-Our mechanisms guarantee the optimal error between each $x_{i}$ and $y_{i}$, which tends to give the smallest $|\hat{\mu}-\mu|$ for any $n$ among GPM.
+Our mechanisms guarantee the optimal error between each $x_{i}$ and $y_{i}$, which in turn leads to the smallest $|\hat{\mu}-\mu|$ among all GPMs under this metric.
 Typically, a mean estimator may also need to be unbiased, i.e.
@@
 This constraint translates to $\mathrm{E}[y_{i}]=\mathrm{E}[\mathcal{M}(x_{i})]=x_{i}$.777 Actually, unbiasedness for numerical data is not as important as for categorical data, as it is for a single data point $x_{i}$.
-For numerical data with infinite possible data points, the probability of having a specific $x_{i}$ is zero.
+When the dataset $\mathcal{X}$ is not concentrated around a single point, an unbiased mechanism may not necessarily provide better performance.
 This is unachievable for same-domain mapping $\mathcal{M}:\mathcal{D}\to\mathcal{D}$ on classical domains, as the endpoints of $\mathcal{D}$ cannot be the mean value (or center) of any distribution over $\mathcal{D}$.
@@
 For any domain $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$, where $\tilde{\mathcal{D}}_{\varepsilon}$ is a variable w.r.t $\varepsilon$, and under absolute error and square error metrics, the optimal piecewise-based mechanism falls into $3$-GPM.
+Hypothesis 5.1 is a natural extension of Hypothesis 3.7, as the output domain $\tilde{\mathcal{D}}_{\varepsilon}$ becomes explicit once $\varepsilon$ is specified.
 Under the $3$-GPM, an unbiased mechanism $\mathcal{M}$ with a variable output domain $\tilde{\mathcal{D}}_{\varepsilon}$ can be analytically derived by incorporating the unbiasedness constraint.
-As a complement to Theorem 3.7, we provide Theorem 5.1 for mean estimation [^8] However, for GPM, unbiasedness does not always result in better data utility for mean estimation, as it comes at the cost of a larger output domain, which leads to higher errors (as shown in Figure 17).
-in the classical domain.
+As a complement to Theorem 3.7, we provide Theorem 5.1 for mean estimation in the classical domain.
 Theorem 5.1.
@@
 Theorems use a concrete domain $\mathcal{D}$ for the sake of clarity.
-We do not give closed-form optimal GPM for distribution estimation on $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$ because $\tilde{\mathcal{D}}_{\varepsilon}$ can not be restricted by constraints as in mean estimation.
+We do not give closed-form optimal GPM for distribution estimation on $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$ because $\tilde{\mathcal{D}}_{\varepsilon}$ can not be easily concretized by constraints as in mean estimation.
 Table 2. Optimal distribution and mean estimation in three domain types under GPM.
@@
 For the best potential of PM and SW, we adapt them to $\mathcal{D}\to\mathcal{D}$ as PM-C and SW-C by linearly compressing their output domain $\tilde{\mathcal{D}}_{\varepsilon}$ to $\mathcal{D}$, i.e.
-transformation invariants, which maintains the privacy level.
-We also compare OGPM’s expected error with non-piecewise-based mechanisms that can be applied to bounded domains:
-  • Variants of the Laplace mechanism: including the staircase mechanism (Geng and Viswanath, 2014), Laplace mechanism with post-processing by truncation (T-Laplace), and the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020), which designs a bounded Laplace-shape distribution.
+by transformation invariants, which maintains the privacy level.
+We also compare OGPM’s expected error with non-piecewise-based mechanisms that can be applied to bounded numerical domains:
+  • Variants of the Laplace mechanism: including the staircase mechanism (Geng and Viswanath, 2014), Laplace mechanism with post-processing by truncation (T-Laplace), and the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020), which redesigns a bounded Laplace-shape distribution.
   • Purkayastha mechanism (Weggenmann and Kerschbaum, 2021): a mechanism for directional data on spheres $\mathbb{S}^{n-1}$. When $n=2$, it is a counterpart of OGPM in the circular domain.
@@
   • whole-domain error: $Err(x)$ values for the whole domain, i.e. for all $x\in\mathcal{D}$.
-  • worst-case error: the largest $Err(x)$ value among the whole domain. PM (Wang et al., 2019) and PTT (Ma et al., 2024) use this error to evaluate data utility.
+  • worst-case error: the largest $Err(x)$ value among the whole domain. PM (Wang et al., 2019) and PTT (Ma et al., 2024) also use this error to evaluate data utility.
 We have proved that the worst-case error of the classical domain is from the endpoints, so this error is actually the error at $x=0$ and $x=1$ for the classical domain.
@@
 T-Laplace preserves the privacy guarantees of the Laplace mechanism while reducing the expected error, particularly for data points near the endpoints or for small $\varepsilon$ values.
-Additionally, the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020) introduces a redesigned bounded Laplace-shaped distribution tailored for bounded domains.999 Appendix B.8 provides details on the expected error of B-Laplace.
+Additionally, the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020) introduces a redesigned bounded Laplace-shaped distribution tailored for bounded domains.888 Appendix B.8 provides details on the expected error of B-Laplace.
 Figure 14. Worst-case error comparison (continued from Figure 13).
@@
 When $n=2$, the sphere $\mathbb{S}^{1}$ corresponds to a circle, making the Purkayastha mechanism a counterpart of OGPM in the circular domain.
-Therefore, we compare them in the circular domain.101010 We omit the comparison with the VMF mechanism also because (i) it has been shown that the Purkayastha mechanism outperforms the VMF mechanism (with the same sensitivity $\Delta_{\measuredangle}=\pi$ for sphere $\mathbb{S}^{1}$, e.g.
+Therefore, we compare them in the circular domain.999 We omit the comparison with the VMF mechanism also because (i) it has been shown that the Purkayastha mechanism outperforms the VMF mechanism (with the same sensitivity $\Delta_{\measuredangle}=\pi$ for sphere $\mathbb{S}^{1}$, e.g.
 Figure 5 and 10 in (Weggenmann and Kerschbaum, 2021)), and (ii) the expected error of the VMF mechanism lacks a closed-form expression (Theorem 17 in (Weggenmann and Kerschbaum, 2021)), making it complex to compute.
@@
 OGPM consistently outperforms other mechanisms in both types of domains, with significantly lower errors.
-In the classical domain, we also compare with OGPM-U, which is designed for unbiased mean estimation, but it has an enlarged output domain, leading to a higher distance between the sensitive and randomized data.
-We can see that such larger output domain results in higher mean estimation errors when $\varepsilon$ is small, but OGPM-U outperforms SW when $\varepsilon\geq 5.5$.
-In the circular domain, OGPM’s mean estimation error is negligible.
-We also observe that SW-C performs better than PM-C in this large-size domain.
+In the classical domain, we also compare with OGPM-U, which is specifically designed for unbiased mean estimation.
+Since the Accelerometer dataset is concentrated around zero, it particularly favors unbiased mechanisms, as their outputs tend to average closely to the true mean.
+We can see that OGPM-U achieves significantly lower errors than OGPM when $\varepsilon$ is small.
+In the circular domain, OGPM is unbiased, having a negligible mean estimation error.
+We also observe that SW-C outperforms PM-C in this large-scale domain.
 Statistically, OGPM’s mean estimation error is $66.2\%$ of PM-C and $55.4\%$ of SW-C in the classical domain.
@@
 Duchi et al.
-randomize any data in $[-1,1)$ to two discrete values $y\in\{-C_{\varepsilon},C_{\varepsilon}\}$ (Duchi et al., 2016), where $C_{\varepsilon}$ is a constant depending on the privacy level $\varepsilon$.
+randomize any data in $[-1,1]$ to two discrete values $y\in\{-C_{\varepsilon},C_{\varepsilon}\}$ (Duchi et al., 2016), where $C_{\varepsilon}$ is a constant depending on the privacy level $\varepsilon$.
 This binary-output mechanism exhibits a large randomization error as the output space is too coarse.
-PM (Wang et al., 2019) extends the binary output to a continuous output in $[-C_{\varepsilon},C_{\varepsilon})$.
+PM (Wang et al., 2019) extends the binary output to a continuous output in $[-C_{\varepsilon},C_{\varepsilon}]$.
 It designs a piecewise-based mechanism to sample the output, which uses different sampling intervals for different $x$, achieving a lower randomization error.
@@
 Another application area is federated learning.
-TPM ensures LDP in the $[0,1)$ domain, which is commonly used as the normalized domain in model training.
-TPM avoids the data clipping required by the noise-adding mechanisms (Abadi et al., 2016; Liu et al., 2023; Fu et al., 2024).
+TPM ensures LDP in the $[0,1]$ domain, which is commonly used as the normalized domain in model training.
+TPM avoids the data clipping required by noise-adding mechanisms (Abadi et al., 2016; Liu et al., 2023; Fu et al., 2024).
 Another typical application is in sensors, where the data is bounded by the sensor’s physical nature.
@@
 Theoretical and experimental evaluations confirm the advantages of our mechanisms over existing mechanisms.
+Acknowledgements.
+We thank the anonymous reviewers and the revision editor for their valuable feedback and guidance, which significantly improved this paper.
+We also acknowledge the use of GPT-4 for language refinement in this paper.
+This research is supported in part by the U.S.
+National Science Foundation under grants CNS-2245689 (CRII) and DGE-2336252, as well as by the 2022 Meta Research Award for Privacy-Enhancing Technologies.
 Appendix A Proofs
@@
 $x$.
-Based on the relationship between $x$ and $[l_{i},r_{i})$, the value of $x$ is split into three cases: (i) $x\in[a,l_{i})$, (ii) $x\in[l_{i},r_{i})$, and (iii) $x\in[r_{i},b)$.
+Specifically, based on the relationship between $x$ and $[l_{i},r_{i})$, the value of $x$ is split into three cases: (i) $x\in[a,l_{i})$, (ii) $x\in[l_{i},r_{i})$, and (iii) $x\in[r_{i},b)$.
 We prove the second derivative of $f_{i}(x)$ w.r.t.
@@
 This statement can be empirically validated by attempting to find counterexamples using larger $m$ than the optimal $m$.
-The source code of our framework provides this validation.
+The source code of our framework provides scripts and results to empirically validate this lemma.
 A.3. Proof of Theorem 3.6
@@
 ∎
-Remark: This proof is the same as finding the optimal $3$-piecewise distribution in domain $[0,1)$. The source code of our framework provides this validation.
+Remark: This proof is the same as finding the optimal $3$-piecewise distribution in domain $[0,1)$. The source code of our framework provides the validation.
 (a) $x=3\pi/4$.
@@
 Focusing on error values allows for more precise comparisons between different mechanisms.
-By excluding post-processing, we can analyze the optimality of the mechanism itself, providing a more fundamental understanding than considering the mechanism combined with specific post-processing.
+By excluding post-processing, we can analyze the optimality of the mechanism itself, which provides a more fundamental understanding than considering the mechanism combined with a specific post-processing.
 Other types of optimality have been explored in the literature, particularly for variants of Laplace mechanisms.
@@
 Formally, if the user’s prior is denoted as $p_{i}$ on the data domain $i\in N$ (i.e.
-a discrete domain) and the user’s post-processing is represented as a remap $z_{i,j}$ that reinterprets the output of the mechanism (on the sensitive value $i$) to $j$, then the utility model is
+a discrete domain) and the user’s post-processing is represented as a remap $z_{i,j}$ that reinterprets the output of the mechanism (on the sensitive value $i$) to $j$, then the utility model is defined as
 $Err(i)=\sum_{i\in N}p_{i}\sum_{j\in N}z_{i,j}\cdot\mathcal{L}(i,j).$
@@
 A mechanism is called universally optimal if, for any prior $p_{i}$, there exists an optimal remap $z_{i,j}$.
-Under this utility model, it was proven that the truncated geometric mechanism (a discretized version of the Laplace mechanism) can achieve universal optimality for count queries111111 This is in the centralized DP setting, where the data curator holds the dataset and uses one mechanism.
+Under this utility model, it was proven that the truncated geometric mechanism (a discretized version of the Laplace mechanism) can achieve universal optimality for count queries101010 This is in the centralized DP setting, where the data curator holds the dataset and uses one mechanism.
 and a legal error metric $\mathcal{L}(i,j)$.
@@
 utility model without prior and post-processing.
-Figure 13 has shown that OGPM generally has a smaller error than the truncated Laplace mechanism, indicating the sub-optimality of the truncated Laplace mechanism in the absence of prior and post-processing.
+Figure 13 has shown that OGPM generally has a smaller error than the truncated Laplace mechanism, especially when the privacy parameter $\varepsilon$ is not small, indicating the sub-optimality of the truncated Laplace mechanism in the absence of using prior and post-processing.
 B.3. Directions for Analytically Proving Optimal $m=3$ (Section 3.2.3)
@@
 transform it into a $3$-staircase distribution while ensuring a decrease in the error, then we can prove that the optimal $m$ is $3$.
-However, this is challenging, as it breaks the probability normalization constraint, requiring adjustments to the probabilities of each piece to satisfy the $\varepsilon$-LDP constraint.
+However, this is challenging, as it breaks the probability normalization constraint (i.e.
+the sum of probabilities is no longer $1$), requiring adjustments to the probabilities of each piece to satisfy the $\varepsilon$-LDP constraint.
 The difficulty lies in ensuring that these adjustments will indeed decrease the error.
@@
 Ideally, these two multi-variable optimization problems could be solved analytically, resulting in two closed-form error expressions w.r.t.
-$x$ and $\varepsilon$, thereby completing the proof for any $x$ and $\varepsilon$.
+$x$ and $\varepsilon$, thereby completing the proof for any $x$ and $\varepsilon$ by showing that the two expressions are equal.
 This direction aligns with our framework.
@@
 Note that $l$ and $r$ should be restricted in $[\tilde{a},\tilde{b})$ when analyzing the first-order derivative.
-B.5. MSE of the Optimal GPM (Section 3.3)
+B.5. MSE of the Optimal GPM (Section 4)
 Denote $p_{\varepsilon}$ and $C$ as the same as the instantiations of $\mathcal{M}$ in Theorem 3.7. The MSE of $\mathcal{M}$ is
@@
 For example, when $x=0$, the MSE of $\mathcal{M}$ is
-$MSE(\mathcal{M}(0),0)=\frac{p_{\varepsilon}}{3}\left(8C^{3}\right)+\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left(1-8C^{3}\right).$
-Setting $\varepsilon=1$ results in $MSE(\mathcal{M}(0),0)=0.22$. As a comparison, SW (Li et al., 2020), which also designed for $\mathcal{D}=[0,1)$, has an MSE of $0.29$ at $x=0$.
+$\mathrm{MSE}[\mathcal{M}(0)]=\frac{p_{\varepsilon}}{3}\left(8C^{3}\right)+\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left(1-8C^{3}\right).$
+Setting $\varepsilon=1$ results in $\mathrm{MSE}[\mathcal{M}(0)]=0.22$ of OGPM. As a comparison, SW (Li et al., 2020), which also designed for $\mathcal{D}=[0,1)$, has an MSE of $0.29$ at $x=0$.
 (a) Privacy parameter $\varepsilon=0.4$.
@@
 B.6. Optimal Assignment of Privacy Parameter and an Example (Section 6)
-The objective function to minimize the total error in 2D polar coordinates is
+The objective function to minimize the given total error in 2D polar coordinates is
 $\min_{\varepsilon_{1},\varepsilon_{2}}Err_{\text{1,wor}}(\varepsilon_{1})+Err_{\text{2,wor}}(\varepsilon_{2}),$
@@
 the classical domain $[0,1)$ is achieved at $x=0$.
-Therefore, $Err_{\text{1,wor}}$ equals to $MSE(\mathcal{M}(0),0)$ calculated before.
+Therefore, $Err_{\text{1,wor}}$ equals to $\mathrm{MSE}[\mathcal{M}(0)]$ calculated before.
 For the circular domain $[0,2\pi)$, the worst-case error $Err_{\text{2,wor}}$ is stated in Theorem 4.4.
 If $\varepsilon=1+2\pi$ and we assign $\varepsilon_{1}$ to the classical domain and $\varepsilon_{2}=\varepsilon-\varepsilon_{1}$ to the circular domain, then the total error is plotted in the right figure.
-The optimal assignment is $\varepsilon_{1}=1.32$ and $\varepsilon_{2}=5.69$.
+In this figure, the optimal assignment is $\varepsilon_{1}=1.32$ and $\varepsilon_{2}=5.69$.
 From the curve of the total error, we can see that $\varepsilon_{2}$ affects the total error more than $\varepsilon_{1}$.
@@
 The B-Laplace mechanism redefines a Laplace-shaped distribution on a bounded domain as the perturbation mechanism.
-For the domain $\mathcal{D}\to\tilde{\mathcal{D}}=[0,1)\to[0,1)$, the B-Laplace mechanism is defined as follows:
+For the data domain $\mathcal{D}\to\tilde{\mathcal{D}}=[0,1)\to[0,1)$, the B-Laplace mechanism is defined as follows:
 Definition B.1 (Bounded Laplace Mechanism, adapted from (Holohan et al., 2020)).